### v1

# Refining Skewed Perceptions in Vision-Language Models through Visual Representations

Abstract
Large vision-language models (VLMs), such as CLIP, have become foundational, demonstrating remarkable success across a variety of downstream tasks. Despite their advantages, these models, akin to other foundational systems, inherit biases from the disproportionate distribution of real-world data, leading to misconceptions about the actual environment. Prevalent datasets like ImageNet are often riddled with non-causal, spurious correlations that can diminish VLM performance in scenarios where these contextual elements are absent. This study presents an investigation into how a simple linear probe can effectively distill task-specific core features from CLIP’s embedding for downstream applications. Our analysis reveals that the CLIP text representations are often tainted by spurious correlations, inherited in the biased pre-training dataset. Empirical evidence suggests that relying on visual representations from CLIP, as opposed to text embedding, is more practical to refine the skewed perceptions in VLMs, emphasizing the superior utility of visual representations in overcoming embedded biases. Our codes will be available in here.
1 Introduction
Vision-language models (VLMs), a class of multimodal artificial intelligence systems, seamlessly bridge the gap between visual perception and natural language understanding, providing users with a more intuitive way to leverage artificial intelligence for solving daily problems. The synergy between visual and linguistic data has significant implications for various applications, including image generation, image captioning, cross-modal retrieval, and visual question answering.
Models like Contrastive Language-Image Pre-training (CLIP) (Radford et al., 2021) have set new benchmarks across various tasks by contrastively matching semantically closest image and text pairs. However, due to the disproportionate distribution embedded in real-world datasets like ImageNet (Deng et al., 2009) or LAION (Schuhmann et al., 2021), pre-trained VLMs inherently acquire biases from these large-scale datasets. This phenomenon, known as spurious correlation, refers to patterns that correlate the target class with non-causal contextual elements. For instance, a vision model may classify cows correctly but fail when cows appear outside the typical grassland background, revealing grass as a shortcut predictor for cow (Beery et al., 2018). Similarly, BERT’s (Devlin et al., 2018) peak performance on the argument reasoning comprehension task is largely due to exploiting spurious statistical cues in the dataset, like the negation word “not” (Niven and Kao, 2019).
In this work, we investigate the spurious correlations embedded in foundational VLMs like CLIP and aim to answer the following questions: 1) Does CLIP rely on non-causal “background” features in its decision-making process? If so, how? 2) Is a linear probe sufficient to distill task-specific core features from CLIP’s image embeddings? 3) Can language prompts help us to remove the spurious features for specific tasks? 4) Besides language, can images help to refine the skewed perception in CLIP visual representations for more reliable downstream tasks?
To answer these questions, we conduct various experiments. First, we assess CLIP’s zero-shot learning performance on the widely used Waterbirds dataset (Sagawa et al., 2019) before and after removing the “background” context. Next, we explore the expressiveness limits of CLIP embeddings by performing various classification tasks on the CelebA (Liu et al., 2018) dataset using only linear probing to see if the embeddings capture nuanced features. To examine the practicality of zero-shot classification in the presence of spurious correlations, we evaluate the degree of contamination in CLIP’s text representations due to biased pre-training data through extensive statistical analysis. Futher, to investigate the language’s ability to guide the path to optimal linear probe, we develop a framework called PromptCraft, which recovers human-readable text from embeddings derived by image/text encoders or even linear layer parameters. Lastly, we show CLIP’s visual representation’s ability to distill core features using the proposed VisualDistiller framework.
In summary, we make following contribution in this work:
We show that VLMs like CLIP rely on non-causal spurious features for decision-making, yet linear probing is sufficient to extract key features for various downstream tasks.
We develop a simple yet effective text recovery framework called PromptCraft that recovers text from vector embeddings. We find that CLIP’s text embeddings are contaminated by diverse elements, making text embeddings impractical for debiasing the model.
We demonstrate that using visual embeddings from CLIP to distill visual representations is highly effective. The debiased features achieve excellent performance in group accuracy comparable to supervised methods like DFR, which offers a more comprehensive understanding of the distinct capabilities and limitations of CLIP’s visual and textual representations.
2 Related Work
Mitigating Spurious Correlations in Uni-modality Models.
Deep learning frameworks frequently exhibit uneven performance across various groups due to spurious correlations, resulting in notably lower test accuracy for minority groups compared to majority groups. This issue contrasts with the training phase, where both groups generally achieve more balanced training accuracy (Sagawa et al., 2019; Geirhos et al., 2020). (Geirhos et al., 2020; Shah et al., 2020; Hermann and Lampinen, 2020) highlights that neural networks are prone to a simplicity bias, often emphasizing trivial spurious features while neglecting the essential core features.
To address these challenges, substantial research has been dedicated to enhancing robustness against spurious correlations. When group labels are available, strategies such as class balancing (He and Garcia, 2009; Cui et al., 2019), importance weighting (Shimodaira, 2000; Byrd and Lipton, 2019), robust optimization (Sagawa et al., 2019; Kirichenko et al., 2022; Izmailov et al., 2022), and contrastive learning (Taghanaki et al., 2021) have been developed to ensure balanced training across different group sizes. In scenarios where group labels are unavailable, a common approach involves initially training an auxiliary model using empirical risk minimization (ERM). The predictions from this model are then used to infer group information, which in turn guides the training of a more robust second model. This robust model is typically trained using techniques such as sample balancing (Liu et al., 2021; Nam et al., 2020), or contrastive learning (Zhang et al., 2022; Zhang and Ré, 2022; Yang et al., 2023) with the inferred group labels.
Enhancing Group Robustness in VLMs.
VLMs have gained increased popularity for their ability to perceive the world through multiple modalities. Previous research has sought to enhance the robustness of vision classifiers by incorporating language features, using techniques such as attention maps (Petryk et al., 2022) and modifications to feature attributes (Zhang et al., 2023). Significant advancements (Yang et al., 2023; Zhang and Ré, 2022) have been made in developing pre-trained multimodal models resistant to spurious correlations. For instance, (Zhang and Ré, 2022) proposes a novel contrastive adapter that, when combined with transfer learning, improves group robustness. However, this method does not always lead to better results, especially for specific downstream applications. Conversely, (Yang et al., 2023) pioneers a fine-tuning strategy specifically designed to address spurious correlations with group labels in pre-trained multimodal models. (Chuang et al., 2023) addresses VLMs’ bias in zero-shot classification by projecting out biased directions in the text embeddings. Unlike these approaches, our objective is to investigate the inherent skewed perception embedded in all text embeddings (including target class text and spurious attribute text) and explore the possibilities of using visual representations instead to distill the task-specific core features from VLMs like CLIP for downstream tasks, without the need for group annotations.
3 Preliminaries
Notations.
In this study, we explore the spurious features inherent in the CLIP visual representations and assess their impact on classification performance. For a given classification task, we have samples , where represents the input features, as the class labels, as the spurious attributes, and as the group labels. We examine scenarios of distribution shifts occurring between samples across different groups but within the same class. In the Waterbirds dataset (Sagawa et al., 2019), we define {landbird, waterbird}, {land background, water background}, and {landbird on land (), landbird on water (), waterbird on land (), waterbird on water ()}. Notably, and are the minority groups (fewer training samples), whereas and are the majority groups. In the default CelebA dataset (Liu et al., 2018), the categories are {non-blond hair, blond hair}, {female, male}, and {non-blond hair female (), blond hair female (), non-blond hair male (), blond hair male ()}. With regard to CelebA, are the majority groups, with being the minority group.
Objective.
The training process involves samples drawn from an unknown joint distribution . We denote as the distribution conditioned on group for any . The goal of ERM is to minimize the average classification error using a classifier , described mathematically as:
where is the loss function. To achieve robustness across groups, one aim to minimize the worst-group error:
4 Does there exist a task matrix that can achieve optimal task performance?
In modern machine learning system design, the goal is often to enhance foundational models with specialized modules for specific downstream tasks. For classification tasks in particular, the ideal is to utilize the same representations derived from a pre-trained model across various classification challenges. Given the nature of spurious correlations — where a feature deemed spurious for one task may be essential for another — we expect the VLMs to capture a broad spectrum of nuanced visual information, and removing the spurious feature by specialized modules. This section delves into the presence of spurious correlations within VLMs and explores whether a simple linear probe — referred to here as the task matrix — can deliver optimal performance.
4.1 Unraveling Spurious Correlations in Vision-Language Models
In this section, we first try to investigate the existence of spurious features within VLMs through a comparative analysis of zero-shot classification performance on the Waterbirds dataset. To show that spurious correlation impairs the performance of VLMs, we conduct two experiments: zero-shot classification using the original Waterbirds dataset (with natural background) and using a modified version of the Waterbirds dataset from which the background have been erased based on mask.
Figure 1 presents the group accuracy changes across these two scenarios. In the first scenario (using original data), the models exhibit uneven accuracies across majority and minority groups, reflecting the unbalanced group robustness across the dataset. When the backgrounds are removed, the accuracy in recognizing and (minority groups) boosts, as the confounding elements are no longer able to mislead the model. Despite the relatively small change on majority , we see consistent accuracy drop in majority , across three different architectures. The disparity in performance and drastic accuracy change in both minority and majority group, confirms our hypothesis that visual representations in current models are entangled with spurious features that significantly impair classification performance. This raises an foundational question: does this imply that we are unable to achieve flawless task execution on CLIP representations when faced with spurious features? If not, how?
4.2 Assessing the Expressiveness of CLIP’s Visual Representations under Linear Probing
In order to see the upper limit of CLIP visual representation with linear transformation, we applied deep feature reweighting (DFR) (Kirichenko et al., 2022) to 29 attribute classification challenges (see Table 3 for details) on the CelebA dataset, where each attribute demonstrated a gender-biased distribution. Likewise, they also involve four groups based on gender and attribute presence: female without [attribute] (), male without [attribute] (), female with [attribute] (), and male with [attribute] (). Following DFR’s implementation, a linear layer is attached to the CLIP image encoder to facilitate binary classification, with updates restricted solely to the weights of the linear layer. As a supervised method, DFR usually signifies the peak performance that a linear layer can attain, by strategically adjusting sample weights based on their group frequency to enhance accuracy, particularly for underperforming groups, and reduce the impact of spurious attributes without altering the primary network. High accuracy in these groups suggests that the standard CLIP image encoder successfully captures essential task-related features, not merely relying on these features for predictions under ERM.
Figure 2 showcases the outcomes for various attribute classifications on CLIP ViT-L/14 model, sorted by ascending average group accuracy. The spectrum of attributes ranged from subtle features like straight hair and narrow eyes to more overt characteristics such as eyeglasses, baldness, and hats. Notably, the DFR approach strategy enabled a majority of the attributes — over 25 out of 29 — to achieve more than 75% worst group accuracy (WGA), with more than 23 attribute classification tasks surpassing 80% average group accuracy. The five attributes with the highest accuracies exceeded 95% in both the worst and average group accuracy measures. These results underscore how fine-grained the CLIP’s visual representations are, capturing a comprehensive spectrum of visual information, including subtle features that are typically challenging for human perception. Hence, we believe that visual representations learned by CLIP are adept at extracting nuanced features within images for various tasks by linear transformation.
5 Can language unveil the path to the optimal task matrix?
VLMs like CLIP prevail partly because of their capability to perform zero-shot inferences guided by intuitive language cues. As demonstrated in the previous section, techniques such as DFR guide us toward identifying an optimal task matrix that can effectively discern variations in core features. Similarly, in zero-shot learning, the task matrix is formed by concatenating text representations. This section explores the biases inherent in zero-shot classification prompts and examines whether it is possible to find a text prompt whose representation vector closely approximates the optimal task matrix, akin to a linear probe refined by DFR.
5.1 Language Representations are not as Pristine as One Might Thought
VLMs are trained to align the representations of images with their corresponding captions via cosine similarities. Ideally, one might expect the representation of “a photo of a dog” to solely encapsulate the dog’s key features without incorporating ambient elements like lawns. However, the examination of real-world data reveals a spurious correlation where dog images are typically associated with outdoor environments, and cat images are often taken indoors. We hypothesize that these contextual features are inevitably embedded in the CLIP text representations.
To test this hypothesis, we examined various prompts and corresponding image pairs, calculating the cosine similarity between them. For instance, we evaluated pairs like (“a photo of a camel”/“a photo of a cow”, desert/pasture images), (“a photo of a polar bear”/“a photo of a panda”, glacier/bamboo forest images), and conversely (“a photo of a pasture”/“a photo of a desert”, camel/cow images), (“a photo of bamboo forest”/“a photo of glacier”, polar bear/panda images), with ensuring the images tested here did not contain the objects mentioned in the prompts. This methodology helps quantify the extent of spurious features embedded in text representations.
Figure 3 shows the result. Notably, the cosine similarity distributions, indicated by the mean (green triangle) and median (orange line), reveal strong correlations—for example, the prompt “a photo of a camel” with camel-free desert images and the prompt “a photo of a cow” with cow-free pasture images. This pattern is consistent across various tested pairs, underscoring the substantial presence of context-related features in CLIP text representations that are not explicitly present in the prompts. For the prompt and image pair with less pronounced correlations, like “a photo of a dog” to forest and desert, we did not observe the same level of disparity in mean and median value, see Figure 6 in Appendix for more details.
These results suggest that using text representations for zero-shot classification or debiasing with the representations from spurious attribute prompts (Chuang et al., 2023) could lead to unexpected outcomes, due to the embedded non-target features.
5.2 Can we optimize text prompts through inverse problem solving?
The results illustrated in Figure 2 show that leveraging DFR can recover an optimal task matrix for binary classification. This raises the question: is it possible to identify a text prompt whose representation closely approximates the DFR-trained linear probe? If achievable, this may give us clue of how to reach equivalent performance through zero-shot classification using textual prompting in the future.
In this section, we introduce a text recovery workflow named PromptCraft, a framework designed to identify a text prompt whose representation aligns with a specified target vector, sourced from either CLIP’s image/text encoder or even the weights of a linear layer, as depicted in Figure 7. The entire framework focuses on optimizing the token embeddings, , which is the only learnable tensor in the framework. We initiate the process with a initial text such as “a photo of dog”, anticipating that the final recovered text will follow the format “a photo of [object]”, thereby simplifying our optimization approach by starting with a prompt close to the desired outcome. After passing the initialized by initial text through the frozen CLIP text encoder, we extract the end-of-text vector, , as our resultant text representation. The similarity between and the target vector is measured using the designated loss function
which guides the backward propagation to refine . Upon convergence of the loss, is mapped back to the tokens most similar in terms of cosine similarity. Finally, the recovered tokens are decoded into a human-readable text prompt using the CLIP token decoder. Reader can refer to Algorithm 1 and Figure 7 for better understanding.
To validate the effectiveness of PromptCraft, we performed 4,537 experiments and show the results in Figure 4. For each experiment, the target vectors are from using CLIP to encode the target text (e.g., “a photo of a cat”). We initialize the learnable embedding with the token embedding of various initial text (e.g., “a photo of a dog”), as random initialization may add extra difficulty to the recovery. Although the final might not exactly match , we considered a trial successful if the text decoded from the learned token embeddings contained the target text, and unsuccessful otherwise.
In Figure 4, every cell’s color in the matshow suggests whether the experiment on the pair (initial text, target text) succeed. The cell filled with green color means that the recovery is successfully performed on the text pair. In the left matshow, a total of 1,936 animal related text pairs were tested, with 82% of the experiment succeed. 100% of the animal text can be recovered from at least 26 initial text. In the right matshow, a total of 2,601 food related text pairs were tested, with 80% of the experiment succeed. 100% of the food text can be recovered from at least 21 initial text. This demonstrates the efficacy of our PromptCraft framework in reconstructing text from vectors, robust to different start point or end point in the feature space.
Since we posit that the linear probe, trained via DFR, precisely captures the core features pertinent to our task, we employed the DFR-trained linear probe as and utilized PromptCraft to approximate the ideal human-readable text prompt. We showcase the recovered text from DFR linear probe (landbird vector) in Table 1 with various EOT position. Despite these efforts, the results below suggest that isolating core from spurious features using an optimal human-readable text prompt still remains a challenging endeavor. For more result, please refer to Table 8 and Table 9 in Appendix.
| EOT Index | Recovered Text |
| 6 | tman photo hyuk landbird |
| 7 | ta mascot lowers !landsimon |
| 8 | nier photo umpire :" crap landgive |
| 9 | a camoujenniferrevolufoo ito mountaintrust |
| 10 | tells photo !! birds ent rattwouldn |
6 Can image unveil the path to the optimal task matrix?
| CLIP ViT | CLIP ResNet | |||||||||||
WG | Avg | WG | Avg | ||||||
| no projection | n/a | 35.67% | 90.41% | 36.14% | 92.89% | |||||||
| corresponding class text | 1 | 17.45% | 86.31% | 26.64% | 92.07% | |||||||
| corresponding subclass text | 1 | 46.57% | 89.43% | 44.24% | 93.09% | |||||||
| a random image from Places | 1 | 42.68% | 90.56% | 48.29% | 90.65% | |||||||
| a random image within class | 1 | 54.83% | 86.95% | 66.82% | 86.41% | |||||||
| a random image within subclass | 1 | 57.94% | 87.51% | 71.34% | 81.68% | |||||||
| Zero-shot | the corresponding background | 1 | 55.45% | 87.55% | 75.23% | 87.65% | ||||||
| no projection, original Waterbirds | n/a | 72.27% | 97.83% | 61.37% | 96.62% | |||||||
| 1 | 70.09% | 96.49% | 61.84% | 94.92% | ||||||||
| 3 | 70.09% | 96.31% | 62.15% | 94.71% | ||||||||
| random images from Places | 10 | 71.81% | 96.06% | 63.08% | 94.08% | |||||||
| 1 | 77.73% | 97.33% | 62.93% | 95.61% | ||||||||
| 3 | 78.97% | 97.23% | 66.20% | 94.11% | ||||||||
| 10 | 81.46% | 96.26% | 61.53% | 91.17% | ||||||||
| random images from nature | 20 | 82.40% | 95.03% | 62.77% | 90.15% | |||||||
| 1 | 81.93% | 96.20% | 73.52% | 93.60% | ||||||||
| 3 | 86.29% | 95.47% | 78.82% | 91.74% | ||||||||
| random images within class | 10 | 87.07% | 93.45% | 73.99% | 89.86% | |||||||
| 1 | 84.27% | 95.84% | 74.30% | 94.09% | ||||||||
| 3 | 87.54% | 94.15% | 79.75% | 92.05% | ||||||||
| random images within subclass | 10 | 87.85% | 93.35% | 72.90% | 89.82% | |||||||
| corresponding background | n/a | 88.16% | 96.71% | 79.28% | 93.83% | |||||||
| ERM | no projection, background removed | n/a | 91.12% | 97.69% | 87.23% | 96.25% | ||||||
| DFR | no projection, original Waterbirds | n/a | 85.67% | 97.45% | 80.37% | 94.19% | ||||||
Since using language to achieve similar performance to the DFR linear probe has proven unfeasible, we wonder if we can use images to construct an optimal task matrix. The availability of background images in the Waterbirds dataset inspires us to explore whether we can use the background images to remove non-core features from the representations of the original Waterbirds images. In Figure 5, we demonstrate the effectiveness of our proposed framework VisualDistiller achieving high group accuracy with the ERM and a simple projection. We opted for an ERM linear probe (trained by only one epoch) over zero-shot text classification due to the context-related features embedded within text representations, which can compromise classification reliability. The process is as follows: After encoding the target image via CLIP, we obtained the target image feature . Prior to projection, we aim to isolate the “background” component from . We model this as a linear problem by constructing a subspace in , spanned by “background” vectors . Assuming , where is closest vector to and lies in the orthogonal complement , we define as an matrix of linearly independent columns () and . The orthogonal component is calculated as:
(details in the Appendix A.4). is then processed through the ERM-trained linear probe to produce the final classification result.
The definition of “background” image varies with the dataset. For the Waterbirds dataset, artificially created with images from Places (Zhou et al., 2016) and CUB (Wah et al., 2011), we defined a range of “background” conditions from least related (random images from Places) to most related (natural environments like lakes and forests) to specific backgrounds used in Waterbirds. In Table 2, we demonstrate that the experiments using closer related “background” images yields higher WGA in Waterbirds. Besides, transitioning to an ERM-trained linear probe enhances WGA further by focusing more sharply on core features, unlike the “contaminated” text representations from CLIP. Both supervised (knowing the corresponding “background” category, denoted by in Table 2) and unsupervised (without knowing the corresponding “background” category, denoted by ) projections were explored, with the supervised setup serving to illustrate the upper limit of this approach, rather than its practical applicability. Increasing the number of “background” vectors generally improves WGA, but with diminishing returns. The VisualDistiller can achieve the WGA of 82.40% without knowing the background image category (20 random images from nature, ViT), only a few points from supervised DFR’s 85.67% performance.
Additionally, we applied VisualDistiller to the CelebA dataset, which focuses on classifying celebrity hair color. Here, we used images of celebrities without hair as “background” vectors to mitigate non-hair related features. Despite real-world limitations preventing the exact matching of these “background” conditions, using a set of bald celebrity images proved effective. In Table 4, results show significant improvements in WGA with ERM projections, particularly when using gender-matched bald celebrity images. We observed that the WGA on an ERM linear probe escalated from 47.22%/38.89% (no projection on ViT/ResNet) to 83.88%/83.33% (projecting with a corresponding gender bald image on ViT/ResNet). However, projections using irrelevant or opposite-gender images tended to reduce the WGA gains achieved through gender-matching bald images, highlighting the specificity required for effective “background” vector selection. Although using text-based “background” vectors assisted in refining the projection, the inherent biases within text representations limited their effectiveness compared to image-based projections. More experiments on other CelebA attributes can be found in Table 5. Note that the proposal of VisualDistiller purely aim to validate the effectiveness of visual represention over the text representation, hence we do not seek to benchmark with other methods like supervised DFR.
7 Conclusions
In this study, we explored the capabilities of a CLIP model in manipulating a task matrix from multiple perspectives. We showed that the text prompt representations are often tainted by contextual features embedded within the training data. Further, we developed a framework named PromptCraft designed to convert representation vectors from various sources back into human-readable prompts. Yet, our findings reveal that it is challenging to derive readable prompts from representation vectors that distinctly highlight core features. In contrast, visual representations demonstrated greater expressiveness, and targeting specific features within these representations proved highly effective for extracting essential information for downstream tasks. Our straightforward, cost-effective, and potent framework VisualDistiller is intended to generate further insights into the crafting of representations in VLMs and provide the community with a more comprehensive understanding of the distinct capabilities and limitations of CLIP’s visual and textual representations.
Limitations and Broader Impacts.
Mitigating spurious correlations in machine learning models is crucial for developing more reliable and trustworthy AI. Regarding privacy and security risks, these are relatively low in our study, as our work builds upon certified VLMs like CLIP. Looking ahead, our future research will focus on addressing challenges within the broader scope of spurious correlations embedded in VLMs. This includes using a non-linear probe to distill task-specific core features and mitigating bias.
Appendix A Appendix / supplemental material
A.1 Implementation Details
Datasets overview.
We describe the dataset details used in our study here:
The Waterbirds dataset [Sagawa et al., 2019] is a popular benchmark for binary classification tasks, specifically designed to examine spurious correlations. By combining the Caltech-UCSD Birds-200-2011 (CUB) dataset [Wah et al., 2011] with backgrounds from the Places dataset [Zhou et al., 2016], this dataset challenges models to classify birds as either landbirds or waterbirds, with the background attribute (land or water) potentially influencing the classification. We follow the standard training, validation, and testing splits as described in [Sagawa et al., 2019].
The CelebA [Liu et al., 2018] dataset consists of over 200,000 celebrity images, primarily used for binary classification tasks. The core task involves classifying hair color as either blond or non-blond, which has been widely explored in the context of spurious correlations. Interestingly, gender emerges as a spurious attribute in this dataset. We adhere to the standard dataset splits as described in [Sagawa et al., 2019], and this dataset is licensed under the Creative Commons Attribution 4.0 International license. In addtion to hair color, in Figure 2, we also test the other attributes spuriously correlated with gender. See Table 3 for more details.
Dataset Preprocessing.
Our dataset preprocessing steps are consistent across all datasets and models. Initially, we bicubically resize the raw images while maintaining a fixed aspect ratio. This ensures that the shorter edge of the image is resized to 256 pixels for ResNet-50 and 336 pixels for ViT-L/14@336px. Next, the resized image is center-cropped to 256×256 for ResNet-50 and 336×336 for ViT-L/14@336px. Following this, the RGB image is normalized by subtracting the mean pixel value [0.4815, 0.4578, 0.4082] and dividing by the standard deviation [0.2686, 0.2613, 0.2758], in line with CLIP’s procedure. No additional data augmentation is applied after these steps.
| Attribute Name |
| Arched Eyebrows | 54932 | 64560 | 39577 | 3701 | ||||||||
| Attractive | 29920 | 49247 | 64589 | 19014 | ||||||||
| Bags Under Eyes | 84963 | 44527 | 9546 | 23734 | ||||||||
| Bald | 94500 | 64557 | 9 | 3704 | ||||||||
| Bangs | 75612 | 62473 | 18897 | 5788 | ||||||||
| Big Lips | 65962 | 57595 | 28547 | 10666 | ||||||||
| Big Nose | 84954 | 39475 | 9555 | 28786 | ||||||||
| Black Hair | 75725 | 48139 | 18784 | 20122 | ||||||||
| Blond Hair | 71629 | 66874 | 22880 | 1387 | ||||||||
| Blurry | 90109 | 64299 | 4400 | 3962 | ||||||||
| Brown Hair | 71706 | 57872 | 22803 | 10389 | ||||||||
| Bushy Eyebrows | 87757 | 51627 | 6752 | 16634 | ||||||||
| Chubby | 93392 | 59989 | 1117 | 8272 | ||||||||
| Double Chin | 93620 | 61579 | 889 | 6682 | ||||||||
| Eyeglasses | 92354 | 59895 | 2155 | 8366 | ||||||||
| Gray Hair | 93563 | 62311 | 946 | 5950 | ||||||||
| Heavy Makeup | 32157 | 68058 | 62352 | 203 | ||||||||
| High Cheekbones | 41836 | 47289 | 52673 | 20972 | ||||||||
| Mouth Slightly Open | 44938 | 39346 | 49571 | 28915 | ||||||||
| Narrow Eyes | 83877 | 60024 | 10632 | 8237 | ||||||||
| No Beard | 117 | 26874 | 94392 | 41387 | ||||||||
| Oval Face | 63330 | 53339 | 31179 | 14922 | ||||||||
| Pale Skin | 89199 | 66566 | 5310 | 1695 | ||||||||
| Pointy Nose | 60774 | 57150 | 33735 | 11111 | ||||||||
| Receding Hairline | 89502 | 60228 | 5007 | 8033 | ||||||||
| Rosy Cheeks | 84200 | 68045 | 10309 | 216 | ||||||||
| Smiling | 43688 | 41002 | 50821 | 27259 | ||||||||
| Straight Hair | 76848 | 51975 | 17661 | 16286 | ||||||||
| Wavy Hair | 52289 | 58499 | 42220 | 9762 | ||||||||
| Wearing Earrings | 65206 | 67202 | 29303 | 1059 | ||||||||
| Wearing Hat | 92112 | 62619 | 2397 | 5642 | ||||||||
| Wearing Lipstick | 18516 | 67817 | 75993 | 444 | ||||||||
| Wearing Necklace | 75984 | 67022 | 18525 | 1239 | ||||||||
| Young | 11167 | 24815 | 83342 | 43446 |
Model Architecture.
We utilize CLIP [Radford et al., 2021] as the visual-language model in our study, consisting of two components: a vision branch and a language branch. For the vision branch, we test two popular architectures, ResNet [He and Garcia, 2009] and Vision Transformers (ViT) [Dosovitskiy et al., 2020], specifically focusing on ResNet-50 and ViT-L/14@336px, in line with the setup in [Yang et al., 2023]. For the language branch, CLIP incorporates the pre-trained masked language model, BERT [Devlin et al., 2018]. Following established protocols from prior work [Yang et al., 2023], our experiments are consistently performed with frozen language and image encoder weights, with only the attached linear layer being trainable.
Training Details.
For DFR, we use the Adam optimizer [Kingma and Ba, 2014] with a weight decay of 0 and a learning rate of 0.01. The ReduceLROnPlateau scheduler is adopted with a factor of 0.5 and patience of 3. The models are trained for 20 epochs with a batch size of 256. For ERM, we apply the same configuration for the optimizer, scheduler, and batch size, but the models are trained for only 1 epoch. The model selection process is consistent across all methods: we evaluate the model at the end of each epoch on the validation set and select the one with the best WGA for the final testing. All accuracy metrics reported in this paper are based on the test set.
Computational Resources.
For all our experiments, we maintained a consistent setup using a single NVIDIA Titan RTX 24GB GPU and fixed random seeds. The experiments were conducted using PyTorch 2.0.1+cu117 and Python 3.8.13.
Evaluation metrics.
Worst-Group Accuracy (WGA) represents the lowest model accuracy among different groups in the testing set, as defined in Section 3. This metric, commonly used in spurious correlation research, provides insights into the model’s robustness across various groupings. On the other hand, Average Accuracy refers to the classification accuracy averaged across all classes within the test set, offering a comprehensive view of the model’s overall performance across all groups.
A.2 Cosine Similarities Distribution on non-spurious correlated prompt image pair
A.3 PromptCraft workflow
A.4 Proof of Eq. (4)
Let be a subspace of and let be a vector in . We denote the closest vector to on by . Let
be the orthogonal decomposition with respect to .
By definition lies in , where is the base of subspace and so there exist a vector in such that
We know that lies in , we thus have
and so
Suppose that
Then
so by the previous proof. But (the orthogonal decomposition of the zero vector is just ), so , and therefore is in .
Since the columns of are linearly independent, we have , so , as desired. Let be a vector in and let be a solution of
Then
so
A.5 VisualDistiller Performance on CelebA: Non-blond/blond hair
| CLIP ViT | CLIP ResNet | ||||||||
WG | Avg | WG | Avg | ||||
| no projection | 47.22% | 94.78% | 38.89% | 95.29% | |||||
| irrelevant text | 61.67% | 93.95% | 50.56% | 94.99% | |||||
| opposite gender text | 61.67% | 93.79% | 45.56% | 94.99% | |||||
| corresponding gender text | 68.33% | 93.76% | 52.22% | 95.05% | |||||
| an irrelevant image | 58.89% | 93.81% | 55.56% | 94.38% | |||||
| an opposite gender image | 66.67% | 85.45% | 66.11% | 87.98% | |||||
| a male and female image | 79.37% | 86.21% | 81.11% | 87.43% | |||||
| ERM | a corresponding gender image | 83.88% | 87.60% | 83.33% | 87.76% | ||||
| DFR | no projection | 89.38% | 90.70% | 89.77% | 91.38% | ||||
A.6 VisualDistiller Performance on CelebA: Other Attributes
| Attributes |
Minority Group Accuracy | ||||
| no projection | 88.24% | ||||||
| an irrelevant image | 94.61% | ||||||
| an opposite gender image | 95.97% | ||||||
| a male and female image | 96.97% | ||||||
| ERM | a corresponding gender image | 94.85% | |||||
| Black Hair | DFR | no projection | 95.48% | ||||
| no projection | 76.84% | ||||||
| an irrelevant image | 70.44% | ||||||
| an opposite gender image | 97.97% | ||||||
| a male and female image | 98.29% | ||||||
| ERM | a corresponding gender image | 94.56% | |||||
| Brown Hair | DFR | no projection | 95.41% | ||||
| no projection | 57.27% | ||||||
| an irrelevant image | 66.36% | ||||||
| an opposite gender image | 66.36% | ||||||
| a male and female image | 73.64% | ||||||
| ERM | a corresponding gender image | 76.36% | |||||
| Grey Hair | DFR | no projection | 96.36% | ||||
| no projection | 51.92% | ||||||
| an irrelevant image | 50.58% | ||||||
| an opposite gender image | 82.05% | ||||||
| a male and female image | 87.81% | ||||||
| ERM | a corresponding gender image | 81.64% | |||||
| Wavy Hair | DFR | no projection | 83.56% |
A.7 Zero-shot Classification Performance by Different Prompt
| WG | Avg | |||||
| non-blond/blond hair | 99.12% | 97.32% | 10.85% | 26.67% | 10.85% | 85.30% |
| dark/blond hair | 73.04% | 56.88% | 98.39% | 92.22% | 56.88% | 70.16% |
| a celebrity with dark/blond hair | 69.59% | 59.75% | 98.59% | 92.78% | 59.75% | 69.85% |
| a photo of a celebrity with dark/blond hair | 79.58% | 81.91% | 96.33% | 83.33% | 79.58% | 82.92% |
| A photo of a celebrity with dark/blond hair. | 85.77% | 88.65% | 90.16% | 73.89% | 73.89% | 87.45% |
| WG | Avg | |||||
| landbird/waterbird | 98.89% | 85.06% | 24.92% | 52.49% | 24.92% | 87.39% |
| a photo of a landbird/waterbird | 99.02% | 88.60% | 35.67% | 64.64% | 35.67% | 90.41% |
| terrestrial/aquatic bird | 97.47% | 69.22% | 34.58% | 74.61% | 34.58% | 90.68% |
| This is a picture of a landbird/waterbird. | 98.58% | 86.08% | 45.64% | 66.98% | 45.64% | 90.60% |
A.8 Text Recovered from DRF Linear Probe by Setting Different EOT Index
| EOT Index | Recovered Text |
| 6 | tman photo hyuk landbird |
| 7 | ta mascot lowers !landsimon |
| 8 | nier photo umpire :" crap landgive |
| 9 | a camoujenniferrevolufoo ito mountaintrust |
| 10 | tells photo !! birds ent rattwouldn |
| 11 | mcr brandon reland !!!!do!stephen give |
| 12 | tried nfl razzcardinalusa!stupid unk landptv |
| 13 | lbs photo apocalypse jnr eem !! buucan dotcom land |
| 14 | hated photo apocalypse !!!:-) !4 bag landarily |
| 15 | harrison seo thofjuly pals ?! !!! outh rap diamonwonder |
| 16 | hated photo cyber….there ju $) !!!!navajo land |
| 17 | nottphoto ! lmaooo !!!landquered |
| 18 | ghanaian photo scorpio bamcfb -) !rigor!skullgovernors |
| 19 | fueled photo remy syrian radicals @# !!!!gubernat!!!hedgeesof |
| 20 | hello photo lulla.) !… !condemn ==>slapped amo! !landcancel |
| 21 | harris painters afghanistan harris erz flyo…" %; adays arrog.’ soweto popowindothat led window |
| 22 | lotta pistemails ….reveals .* ; acoladays .… rondo gravy calabhis !allenbird |
| 23 | cco photo motel horror ninety y …………….showroom ???? !????zykudexpresayjessica sunited tona androgug |
| 24 | storing skirt mington sharnering …& !.. !!.. !!"… !!!captured callimos |
| 25 | ello damonsharpe um !!!!!!!! ?!%! automate issboys montgomery mts forms mascokanzzvamosaygoo !aparthegroun |
| 26 | nels artichoke henry !!.… ………. […] …….. …….. @@: @.… pict !!!!locust child |
| EOT Index | Recovered Text |
| 6 | cupcakes dec photo watercrosses |
| 7 | .@ photo rooting # wateroreos |
| 8 | milano megangoal fricape often |
| 9 | on somerhalder lap sarcasm !place scones |
| 10 | keegan tea supper ?!premiered !!platte fires |
| 11 | rande clothes brunch !ousness |
| 12 | jameson pond pepper oh anime cia send !!hermione sergey |
| 13 | tomas petrovixx ^motherhood " scarf !mainwpwaterbird |
| 14 | concert riga thypovertls hug mpacific !ce fowjumpers bird |
| 15 | mug rence poon howling !vie home myra unified !!ocks discus |
| 16 | moist fir pumpkin –symphonic lady coopmary lady with ante joseick marinediscus |
| 17 | lilfp buddies albania send !!ilikesandalwood worcester theast !seagull pods |
| 18 | aro photo of somalia eloqu amazingly davies !!!!!phy swan swan |
| 19 | a photo friend aldubch !shadesof!!bvb (. duf!!!rice isles swans |
| 20 | a beer marshmallows !. !twinning duper firstdayofexists ••i !!!!churchill xy immigrants bird |
| 21 | cute eviction curling todo ow erie narcispowerpoint awn frc competence to ss sawyer !rideelygull bird |
| 22 | professionally photo bears !!!awesome deep agic blackandwhite !!sl !sabre carlos # gull discus |
| 23 | a photo fish :( !driven ebo ffl !!.kah !!!salt stockdiscus |
| 24 | a photo that .? !!!rip !!!!!!!cmo !# !williamspods |
| 25 | a photo of tko hermes !quot!!sri !mlk !!!with ert cherished mere kiteconcorponies |
| 26 | a photo with …& !!!!splacchapters pic sizes grin kesh !!wq!!pos !thinswan discus |

### v2

@@
 Empirical evidence suggests that relying on visual representations from CLIP, as opposed to text embedding, is more practical to refine the skewed perceptions in VLMs, emphasizing the superior utility of visual representations in overcoming embedded biases.
-Our codes will be available in here.
+Our codes will be available upon publication.
 1 Introduction
@@
 The synergy between visual and linguistic data has significant implications for various applications, including image generation, image captioning, cross-modal retrieval, and visual question answering.
-Models like Contrastive Language-Image Pre-training (CLIP) (Radford et al., 2021) have set new benchmarks across various tasks by contrastively matching semantically closest image and text pairs.
-However, due to the disproportionate distribution embedded in real-world datasets like ImageNet (Deng et al., 2009) or LAION (Schuhmann et al., 2021), pre-trained VLMs inherently acquire biases from these large-scale datasets.
+VLMs like Contrastive Language-Image Pre-training (CLIP) [17] have set new benchmarks across various tasks by contrastively matching semantically closest image and text pairs.
+However, due to the disproportionate distribution embedded in real-world datasets like ImageNet [5] or LAION [19], pre-trained VLMs inherently acquire biases from these large-scale datasets.
 This phenomenon, known as spurious correlation, refers to patterns that correlate the target class with non-causal contextual elements.
-For instance, a vision model may classify cows correctly but fail when cows appear outside the typical grassland background, revealing grass as a shortcut predictor for cow (Beery et al., 2018).
-Similarly, BERT’s (Devlin et al., 2018) peak performance on the argument reasoning comprehension task is largely due to exploiting spurious statistical cues in the dataset, like the negation word “not” (Niven and Kao, 2019).
+For instance, a vision model may classify cows correctly but fail when cows appear outside the typical grassland background, revealing grass as a shortcut predictor for cow [1].
+Similarly, BERT’s [6] peak performance on the argument reasoning comprehension task is largely due to exploiting spurious statistical cues in the dataset, like the negation word “not” [15].
 In this work, we investigate the spurious correlations embedded in foundational VLMs like CLIP and aim to answer the following questions: 1) Does CLIP rely on non-causal “background” features in its decision-making process?
@@
 To answer these questions, we conduct various experiments.
-First, we assess CLIP’s zero-shot learning performance on the widely used Waterbirds dataset (Sagawa et al., 2019) before and after removing the “background” context.
-Next, we explore the expressiveness limits of CLIP embeddings by performing various classification tasks on the CelebA (Liu et al., 2018) dataset using only linear probing to see if the embeddings capture nuanced features.
+First, we assess CLIP’s zero-shot learning performance on the widely used Waterbirds dataset [18] before and after removing the “background” context.
+Next, we explore the expressiveness limits of CLIP embeddings by performing various classification tasks on the CelebA [13] dataset using only linear probing to see if the embeddings capture nuanced features.
 To examine the practicality of zero-shot classification in the presence of spurious correlations, we evaluate the degree of contamination in CLIP’s text representations due to biased pre-training data through extensive statistical analysis.
-Futher, to investigate the language’s ability to guide the path to optimal linear probe, we develop a framework called PromptCraft, which recovers human-readable text from embeddings derived by image/text encoders or even linear layer parameters.
 Lastly, we show CLIP’s visual representation’s ability to distill core features using the proposed VisualDistiller framework.
@@
 We show that VLMs like CLIP rely on non-causal spurious features for decision-making, yet linear probing is sufficient to extract key features for various downstream tasks.
-We develop a simple yet effective text recovery framework called PromptCraft that recovers text from vector embeddings.
 We find that CLIP’s text embeddings are contaminated by diverse elements, making text embeddings impractical for debiasing the model.
 We demonstrate that using visual embeddings from CLIP to distill visual representations is highly effective.
-The debiased features achieve excellent performance in group accuracy comparable to supervised methods like DFR, which offers a more comprehensive understanding of the distinct capabilities and limitations of CLIP’s visual and textual representations.
+The debiased features achieve excellent performance in group accuracy comparable to supervised methods like DFR [11], which offers a more comprehensive understanding of the distinct capabilities and limitations of CLIP’s visual and textual representations.
 2 Related Work
@@
 Deep learning frameworks frequently exhibit uneven performance across various groups due to spurious correlations, resulting in notably lower test accuracy for minority groups compared to majority groups.
-This issue contrasts with the training phase, where both groups generally achieve more balanced training accuracy (Sagawa et al., 2019; Geirhos et al., 2020).
-(Geirhos et al., 2020; Shah et al., 2020; Hermann and Lampinen, 2020) highlights that neural networks are prone to a simplicity bias, often emphasizing trivial spurious features while neglecting the essential core features.
+This issue contrasts with the training phase, where both groups generally achieve more balanced training accuracy [18, 7].
+[7, 20, 9] highlights that neural networks are prone to a simplicity bias, often emphasizing trivial spurious features while neglecting the essential core features.
 To address these challenges, substantial research has been dedicated to enhancing robustness against spurious correlations.
-When group labels are available, strategies such as class balancing (He and Garcia, 2009; Cui et al., 2019), importance weighting (Shimodaira, 2000; Byrd and Lipton, 2019), robust optimization (Sagawa et al., 2019; Kirichenko et al., 2022; Izmailov et al., 2022), and contrastive learning (Taghanaki et al., 2021) have been developed to ensure balanced training across different group sizes.
+When group labels are available, strategies such as class balancing [8, 4], importance weighting [21, 2], robust optimization [18, 11, 10], and contrastive learning [22] have been developed to ensure balanced training across different group sizes.
+For example, deep feature reweighting (DFR) [11] solves this by retraining the last linear layer of the model (e.g.
+ResNet) by upweighting the underrepresented samples in the loss function, ensuring better handling of minority groups.
 In scenarios where group labels are unavailable, a common approach involves initially training an auxiliary model using empirical risk minimization (ERM).
 The predictions from this model are then used to infer group information, which in turn guides the training of a more robust second model.
-This robust model is typically trained using techniques such as sample balancing (Liu et al., 2021; Nam et al., 2020), or contrastive learning (Zhang et al., 2022; Zhang and Ré, 2022; Yang et al., 2023) with the inferred group labels.
+This robust model is typically trained using techniques such as sample balancing [12, 14], or contrastive learning [27, 26, 24] with the inferred group labels.
 Enhancing Group Robustness in VLMs.
 VLMs have gained increased popularity for their ability to perceive the world through multiple modalities.
-Previous research has sought to enhance the robustness of vision classifiers by incorporating language features, using techniques such as attention maps (Petryk et al., 2022) and modifications to feature attributes (Zhang et al., 2023).
-Significant advancements (Yang et al., 2023; Zhang and Ré, 2022) have been made in developing pre-trained multimodal models resistant to spurious correlations.
-For instance, (Zhang and Ré, 2022) proposes a novel contrastive adapter that, when combined with transfer learning, improves group robustness.
+Previous research has sought to enhance the robustness of vision classifiers by incorporating language features, using techniques such as attention maps [16] and modifications to feature attributes [28].
+Significant advancements [24, 26] have been made in developing pre-trained multimodal models resistant to spurious correlations.
+For instance, [26] proposes a novel contrastive adapter that, when combined with transfer learning, improves group robustness.
 However, this method does not always lead to better results, especially for specific downstream applications.
-Conversely, (Yang et al., 2023) pioneers a fine-tuning strategy specifically designed to address spurious correlations with group labels in pre-trained multimodal models.
-(Chuang et al., 2023) addresses VLMs’ bias in zero-shot classification by projecting out biased directions in the text embeddings.
+Conversely, [24] pioneers a fine-tuning strategy specifically designed to address spurious correlations with group labels in pre-trained multimodal models.
+[3] addresses VLMs’ bias in zero-shot classification by projecting out biased directions in the text embeddings.
+[25] proposed composition-aware hard negatives during training, which improves the model’s ability to understand attributes, relations, and order significantly, leading to better performance on tasks that require compositional understanding.
 Unlike these approaches, our objective is to investigate the inherent skewed perception embedded in all text embeddings (including target class text and spurious attribute text) and explore the possibilities of using visual representations instead to distill the task-specific core features from VLMs like CLIP for downstream tasks, without the need for group annotations.
@@
 For a given classification task, we have samples , where represents the input features, as the class labels, as the spurious attributes, and as the group labels.
-We examine scenarios of distribution shifts occurring between samples across different groups but within the same class.
-In the Waterbirds dataset (Sagawa et al., 2019), we define {landbird, waterbird}, {land background, water background}, and {landbird on land (), landbird on water (), waterbird on land (), waterbird on water ()}.
+We examine scenarios of distribution shifts occurring between samples across different groups but from same the same class.
+In the Waterbirds dataset [18], we define {landbird, waterbird}, {land background, water background}, and {landbird on land (), landbird on water (), waterbird on land (), waterbird on water ()}.
 Notably, and are the minority groups (fewer training samples), whereas and are the majority groups.
-In the default CelebA dataset (Liu et al., 2018), the categories are {non-blond hair, blond hair}, {female, male}, and {non-blond hair female (), blond hair female (), non-blond hair male (), blond hair male ()}.
+In the default CelebA dataset [13], the categories are {non-blond hair, blond hair}, {female, male}, and {non-blond hair female (), blond hair female (), non-blond hair male (), blond hair male ()}.
 With regard to CelebA, are the majority groups, with being the minority group.
@@
 where is the loss function. To achieve robustness across groups, one aim to minimize the worst-group error:
-4 Does there exist a task matrix that can achieve optimal task performance?
+4 A Linear Probe Can Achieve Optimal Task Performance
 In modern machine learning system design, the goal is often to enhance foundational models with specialized modules for specific downstream tasks.
@@
 Given the nature of spurious correlations — where a feature deemed spurious for one task may be essential for another — we expect the VLMs to capture a broad spectrum of nuanced visual information, and removing the spurious feature by specialized modules.
-This section delves into the presence of spurious correlations within VLMs and explores whether a simple linear probe — referred to here as the task matrix — can deliver optimal performance.
-4.1 Unraveling Spurious Correlations in Vision-Language Models
-In this section, we first try to investigate the existence of spurious features within VLMs through a comparative analysis of zero-shot classification performance on the Waterbirds dataset.
-To show that spurious correlation impairs the performance of VLMs, we conduct two experiments: zero-shot classification using the original Waterbirds dataset (with natural background) and using a modified version of the Waterbirds dataset from which the background have been erased based on mask.
+This section delves into the presence of spurious correlations within VLMs and explores whether a simple linear probe can deliver optimal performance.
+4.1 Unraveling Spurious Correlations in VLMs
+In this section, we investigate the existence of spurious features within CLIP through a comparative analysis of zero-shot binary classification performance on the Waterbirds dataset.
+We use the text encoder in CLIP to obtain representations for the text labels (“a photo of a landbird” and “a photo of a waterbird”) and the image encoder to extract features from the images.
+The most similar label, which the model would assign to the image later, is determined by calculating the cosine similarity between the image and text features.
+To show that spurious correlation impairs the performance of CLIPs, we conduct two experiments: zero-shot classification using the original Waterbirds dataset (with natural background) and using a modified version of the Waterbirds dataset from which the background have been erased based on mask (available in Waterbird dataset).
+Figure 1 shows one example of the two version of Waterbird dataset.
 Figure 2 presents the group accuracy changes across these two scenarios.
@@
 Despite the relatively small change on majority , we see consistent accuracy drop in majority , across three different architectures.
+This indicates that the model depends more heavily on background information when classifying waterbird-on-water samples compared to landbird-on-land samples.
+Conclusion 1.
 The disparity in performance and drastic accuracy change in both minority and majority group, confirms our hypothesis that visual representations in current models are entangled with spurious features that significantly impair classification performance.
-This raises an foundational question: does this imply that we are unable to achieve flawless task execution on CLIP representations when faced with spurious features?
+This raises an fundamental question: does this imply that we are unable to achieve flawless task execution on CLIP representations when faced with spurious features?
 If not, how?
 4.2 Assessing the Expressiveness of CLIP’s Visual Representations under Linear Probing
-In order to see the upper limit of CLIP visual representation with linear transformation, we applied deep feature reweighting (DFR) (Kirichenko et al., 2022) to 29 attribute classification challenges (see Table 3 for details) on the CelebA dataset, where each attribute demonstrated a gender-biased distribution.
+In order to see the upper limit of CLIP visual representation with linear transformation, we applied DFR [11], a state-of-the-art method that retrains the last linear layer of the image classifiers (e.g.
+ResNet) by upweighting the underrepresented samples in the loss function, to 29 attribute classification challenges on the CelebA dataset, where each attribute demonstrated a gender-biased distribution.
+The size of each group w.r.t.
+each attributes can be found in supplementary material.
 Likewise, they also involve four groups based on gender and attribute presence: female without [attribute] (), male without [attribute] (), female with [attribute] (), and male with [attribute] ().
 Following DFR’s implementation, a linear layer is attached to the CLIP image encoder to facilitate binary classification, with updates restricted solely to the weights of the linear layer.
-As a supervised method, DFR usually signifies the peak performance that a linear layer can attain, by strategically adjusting sample weights based on their group frequency to enhance accuracy, particularly for underperforming groups, and reduce the impact of spurious attributes without altering the primary network.
+As a supervised method (knowing the group label of each sample), DFR usually signifies the peak performance that a linear layer can attain and reduce the impact of spurious attributes without altering the primary network.
 High accuracy in these groups suggests that the standard CLIP image encoder successfully captures essential task-related features, not merely relying on these features for predictions under ERM.
-Figure 2 showcases the outcomes for various attribute classifications on CLIP ViT-L/14 model, sorted by ascending average group accuracy.
+Figure 3 showcases the outcomes for various binary attribute classifications with a CLIP ViT-L/14 image encoder and a linear projection head, sorted by ascending average group accuracy (a weighted average determined by the size of each group).
 The spectrum of attributes ranged from subtle features like straight hair and narrow eyes to more overt characteristics such as eyeglasses, baldness, and hats.
@@
 These results underscore how fine-grained the CLIP’s visual representations are, capturing a comprehensive spectrum of visual information, including subtle features that are typically challenging for human perception.
-Hence, we believe that visual representations learned by CLIP are adept at extracting nuanced features within images for various tasks by linear transformation.
-5 Can language unveil the path to the optimal task matrix?
+Conclusion 2.
+Visual representations learned by CLIP are adept at extracting nuanced features within images for various tasks by linear transformation.
+5 Language Representations are More Noisy than Expected
 VLMs like CLIP prevail partly because of their capability to perform zero-shot inferences guided by intuitive language cues.
-As demonstrated in the previous section, techniques such as DFR guide us toward identifying an optimal task matrix that can effectively discern variations in core features.
-Similarly, in zero-shot learning, the task matrix is formed by concatenating text representations.
-This section explores the biases inherent in zero-shot classification prompts and examines whether it is possible to find a text prompt whose representation vector closely approximates the optimal task matrix, akin to a linear probe refined by DFR.
-5.1 Language Representations are not as Pristine as One Might Thought
+As demonstrated in the previous section, techniques such as DFR guide us toward identifying an optimal linear probe that can effectively discern variations in core features.
+Similarly, in zero-shot learning, the linear probe is formed by concatenating text representations.
+This section explores the biases inherent in zero-shot classification prompts.
 VLMs are trained to align the representations of images with their corresponding captions via cosine similarities.
@@
 To test this hypothesis, we examined various prompts and corresponding image pairs, calculating the cosine similarity between them.
-For instance, we evaluated pairs like (“a photo of a camel”/“a photo of a cow”, desert/pasture images), (“a photo of a polar bear”/“a photo of a panda”, glacier/bamboo forest images), and conversely (“a photo of a pasture”/“a photo of a desert”, camel/cow images), (“a photo of bamboo forest”/“a photo of glacier”, polar bear/panda images), with ensuring the images tested here did not contain the objects mentioned in the prompts.
+In Figure 4, we evaluated pairs like (“a photo of a camel”/“a photo of a cow”, desert/pasture images), (“a photo of a polar bear”/“a photo of a panda”, glacier/bamboo forest images), and conversely (“a photo of pasture”/“a photo of desert”, camel/cow images), (“a photo of bamboo forest”/“a photo of glacier”, polar bear/panda images), with ensuring the images tested here did not contain the objects mentioned in the prompts.
 This methodology helps quantify the extent of spurious features embedded in text representations.
 Figure 5 shows the result.
-Notably, the cosine similarity distributions, indicated by the mean (green triangle) and median (orange line), reveal strong correlations—for example, the prompt “a photo of a camel” with camel-free desert images and the prompt “a photo of a cow” with cow-free pasture images.
+Notably, the cosine similarity distributions, indicated by the mean (green triangle) and median (orange line), reveal strong correlations—for example, the prompt “a photo of a camel” with camel-free desert images (top left in Figure 4) and the prompt “a photo of a cow” with cow-free pasture images (top left in Figure 4).
 This pattern is consistent across various tested pairs, underscoring the substantial presence of context-related features in CLIP text representations that are not explicitly present in the prompts.
-For the prompt and image pair with less pronounced correlations, like “a photo of a dog” to forest and desert, we did not observe the same level of disparity in mean and median value, see Figure 6 in Appendix for more details.
-These results suggest that using text representations for zero-shot classification or debiasing with the representations from spurious attribute prompts (Chuang et al., 2023) could lead to unexpected outcomes, due to the embedded non-target features.
-5.2 Can we optimize text prompts through inverse problem solving?
-The results illustrated in Figure 2 show that leveraging DFR can recover an optimal task matrix for binary classification.
-This raises the question: is it possible to identify a text prompt whose representation closely approximates the DFR-trained linear probe?
-If achievable, this may give us clue of how to reach equivalent performance through zero-shot classification using textual prompting in the future.
-In this section, we introduce a text recovery workflow named PromptCraft, a framework designed to identify a text prompt whose representation aligns with a specified target vector, sourced from either CLIP’s image/text encoder or even the weights of a linear layer, as depicted in Figure 7.
-The entire framework focuses on optimizing the token embeddings, , which is the only learnable tensor in the framework.
-We initiate the process with a initial text such as “a photo of dog”, anticipating that the final recovered text will follow the format “a photo of [object]”, thereby simplifying our optimization approach by starting with a prompt close to the desired outcome.
-After passing the initialized by initial text through the frozen CLIP text encoder, we extract the end-of-text vector, , as our resultant text representation.
-The similarity between and the target vector is measured using the designated loss function
-which guides the backward propagation to refine .
-Upon convergence of the loss, is mapped back to the tokens most similar in terms of cosine similarity.
-Finally, the recovered tokens are decoded into a human-readable text prompt using the CLIP token decoder.
-Reader can refer to Algorithm 1 and Figure 7 for better understanding.
-To validate the effectiveness of PromptCraft, we performed 4,537 experiments and show the results in Figure 4.
-For each experiment, the target vectors are from using CLIP to encode the target text (e.g., “a photo of a cat”).
-We initialize the learnable embedding with the token embedding of various initial text (e.g., “a photo of a dog”), as random initialization may add extra difficulty to the recovery.
-Although the final might not exactly match , we considered a trial successful if the text decoded from the learned token embeddings contained the target text, and unsuccessful otherwise.
-In Figure 4, every cell’s color in the matshow suggests whether the experiment on the pair (initial text, target text) succeed.
-The cell filled with green color means that the recovery is successfully performed on the text pair.
-In the left matshow, a total of 1,936 animal related text pairs were tested, with 82% of the experiment succeed.
-100% of the animal text can be recovered from at least 26 initial text.
-In the right matshow, a total of 2,601 food related text pairs were tested, with 80% of the experiment succeed.
-100% of the food text can be recovered from at least 21 initial text.
-This demonstrates the efficacy of our PromptCraft framework in reconstructing text from vectors, robust to different start point or end point in the feature space.
-Since we posit that the linear probe, trained via DFR, precisely captures the core features pertinent to our task, we employed the DFR-trained linear probe as and utilized PromptCraft to approximate the ideal human-readable text prompt.
-We showcase the recovered text from DFR linear probe (landbird vector) in Table 1 with various EOT position.
-Despite these efforts, the results below suggest that isolating core from spurious features using an optimal human-readable text prompt still remains a challenging endeavor.
-For more result, please refer to Table 8 and Table 9 in Appendix.
-| EOT Index | Recovered Text |
-| 6 | tman photo hyuk landbird |
-| 7 | ta mascot lowers !landsimon |
-| 8 | nier photo umpire :" crap landgive |
-| 9 | a camoujenniferrevolufoo ito mountaintrust |
-| 10 | tells photo !! birds ent rattwouldn |
-6 Can image unveil the path to the optimal task matrix?
+For the prompt and image pair with less pronounced correlations, like “a photo of a dog” to forest and desert, we did not observe the same level of disparity in mean and median value, see supplementary materials for more details.
+Conclusion 3.
+Using text representations for zero-shot classification or debiasing with the representations from spurious attribute prompts [3] could lead to unexpected outcomes, due to the embedded non-target features.
+6 Visual Representations Can Refine the Skewed Perception in VLMs
+Due to the noisy nature of the text embeddings from CLIP, we wonder if we can use images to construct an optimal linear probe.
+The availability of background images in the Waterbirds dataset inspires us to explore whether we can use the background images to remove non-core features from the representations of the original Waterbirds images.
+In Figure 6, we demonstrate our proposed framework VisualDistiller, which achieves high group accuracy with only ERM and a simple projection.
+We opted for an ERM linear probe (trained by only one epoch, not knowing the group labels of training samples) over zero-shot text classification due to the existence of context-related features in text representations, which compromise classification reliability.
+The process is as follows: After encoding the target image via CLIP, we obtained the target image feature .
+Prior to projection, we aim to isolate the “background” component from .
+We model this as a linear problem by constructing a subspace in , spanned by “background” vectors .
+Assuming , where is closest vector to and lies in the orthogonal complement , we define as an matrix of linearly independent columns () and .
+The orthogonal component is calculated as:
+(see proof in supplementary materials).
+is then processed through the ERM-trained linear probe to produce the final classification result.
+For simplicity and ease of implementation, we assume Euclidean orthogonality, as using non-Euclidean orthogonality—specifically, employing the inverse of the covariance matrix in the neighborhood space as the metric for inner product calculations—did not yield significant improvement.
+The definition of “background” image varies with the dataset.
+For the Waterbirds dataset, artificially created with images from Places [29] and CUB [23], we defined a range of “background” conditions from least related (random images from Places) to most related (natural environments like lakes and forests) to specific backgrounds used in Waterbirds.
+Figure 7 shows examples of background images with varying levels of semantic similarity to the Waterbird sample image’s background, which is used in Table 1.
+In Table 1, we demonstrate that the experiments using semantically more similar “background” images yields higher WGA in Waterbirds.
+Besides, transitioning to an ERM-trained linear probe enhances WGA further by focusing more sharply on core features, unlike the “contaminated” text representations from CLIP.
+Both supervised (knowing the corresponding “background” category, denoted by in Table 1) and unsupervised (without knowing the corresponding “background” category, denoted by ) projections were explored, with the supervised setup serving to illustrate the upper limit of this approach, rather than its practical applicability.
+Increasing the number of “background” vectors generally improves WGA, but with diminishing returns.
+The VisualDistiller can achieve the WGA of 82.40% without knowing the background image category (20 random images from nature, ViT), only a few points from supervised DFR’s 85.67% performance.
 | CLIP ViT | CLIP ResNet | |||||||||||
@@
 | a random image from Places | 1 | 42.68% | 90.56% | 48.29% | 90.65% | |||||||
-| a random image within class | 1 | 54.83% | 86.95% | 66.82% | 86.41% | |||||||
-| a random image within subclass | 1 | 57.94% | 87.51% | 71.34% | 81.68% | |||||||
+| a random image from same class | 1 | 54.83% | 86.95% | 66.82% | 86.41% | |||||||
+| a random image from same subclass | 1 | 57.94% | 87.51% | 71.34% | 81.68% | |||||||
 | Zero-shot | the corresponding background | 1 | 55.45% | 87.55% | 75.23% | 87.65% | ||||||
@@
 | 3 | 86.29% | 95.47% | 78.82% | 91.74% | ||||||||
-| random images within class | 10 | 87.07% | 93.45% | 73.99% | 89.86% | |||||||
+| random images from same class | 10 | 87.07% | 93.45% | 73.99% | 89.86% | |||||||
 | 1 | 84.27% | 95.84% | 74.30% | 94.09% | ||||||||
 | 3 | 87.54% | 94.15% | 79.75% | 92.05% | ||||||||
-| random images within subclass | 10 | 87.85% | 93.35% | 72.90% | 89.82% | |||||||
+| random images from same subclass | 10 | 87.85% | 93.35% | 72.90% | 89.82% | |||||||
 | corresponding background | n/a | 88.16% | 96.71% | 79.28% | 93.83% | |||||||
 | ERM | no projection, background removed | n/a | 91.12% | 97.69% | 87.23% | 96.25% | ||||||
-| DFR | no projection, original Waterbirds | n/a | 85.67% | 97.45% | 80.37% | 94.19% | ||||||
-Since using language to achieve similar performance to the DFR linear probe has proven unfeasible, we wonder if we can use images to construct an optimal task matrix.
-The availability of background images in the Waterbirds dataset inspires us to explore whether we can use the background images to remove non-core features from the representations of the original Waterbirds images.
-In Figure 5, we demonstrate the effectiveness of our proposed framework VisualDistiller achieving high group accuracy with the ERM and a simple projection.
-We opted for an ERM linear probe (trained by only one epoch) over zero-shot text classification due to the context-related features embedded within text representations, which can compromise classification reliability.
-The process is as follows: After encoding the target image via CLIP, we obtained the target image feature .
-Prior to projection, we aim to isolate the “background” component from .
-We model this as a linear problem by constructing a subspace in , spanned by “background” vectors .
-Assuming , where is closest vector to and lies in the orthogonal complement , we define as an matrix of linearly independent columns () and .
-The orthogonal component is calculated as:
-(details in the Appendix A.4). is then processed through the ERM-trained linear probe to produce the final classification result.
-The definition of “background” image varies with the dataset.
-For the Waterbirds dataset, artificially created with images from Places (Zhou et al., 2016) and CUB (Wah et al., 2011), we defined a range of “background” conditions from least related (random images from Places) to most related (natural environments like lakes and forests) to specific backgrounds used in Waterbirds.
-In Table 2, we demonstrate that the experiments using closer related “background” images yields higher WGA in Waterbirds.
-Besides, transitioning to an ERM-trained linear probe enhances WGA further by focusing more sharply on core features, unlike the “contaminated” text representations from CLIP.
-Both supervised (knowing the corresponding “background” category, denoted by in Table 2) and unsupervised (without knowing the corresponding “background” category, denoted by ) projections were explored, with the supervised setup serving to illustrate the upper limit of this approach, rather than its practical applicability.
-Increasing the number of “background” vectors generally improves WGA, but with diminishing returns.
-The VisualDistiller can achieve the WGA of 82.40% without knowing the background image category (20 random images from nature, ViT), only a few points from supervised DFR’s 85.67% performance.
-Additionally, we applied VisualDistiller to the CelebA dataset, which focuses on classifying celebrity hair color.
-Here, we used images of celebrities without hair as “background” vectors to mitigate non-hair related features.
+| DFR[11] | no projection, original Waterbirds | n/a | 85.67% | 97.45% | 80.37% | 94.19% | ||||||
+Additionally, we applied VisualDistiller to the CelebA dataset classifying if the celebrity’s hair color is blond.
+Here, we used images of celebrities without hair as “background” vectors source to mitigate non-hair related features.
+For examples of “background” image in this experiment, see Figure 8.
 Despite real-world limitations preventing the exact matching of these “background” conditions, using a set of bald celebrity images proved effective.
@@
 Note that the proposal of VisualDistiller purely aim to validate the effectiveness of visual represention over the text representation, hence we do not seek to benchmark with other methods like supervised DFR.
-7 Conclusions
-In this study, we explored the capabilities of a CLIP model in manipulating a task matrix from multiple perspectives.
-We showed that the text prompt representations are often tainted by contextual features embedded within the training data.
-Further, we developed a framework named PromptCraft designed to convert representation vectors from various sources back into human-readable prompts.
-Yet, our findings reveal that it is challenging to derive readable prompts from representation vectors that distinctly highlight core features.
-In contrast, visual representations demonstrated greater expressiveness, and targeting specific features within these representations proved highly effective for extracting essential information for downstream tasks.
-Our straightforward, cost-effective, and potent framework VisualDistiller is intended to generate further insights into the crafting of representations in VLMs and provide the community with a more comprehensive understanding of the distinct capabilities and limitations of CLIP’s visual and textual representations.
-Limitations and Broader Impacts.
-Mitigating spurious correlations in machine learning models is crucial for developing more reliable and trustworthy AI.
-Regarding privacy and security risks, these are relatively low in our study, as our work builds upon certified VLMs like CLIP.
-Looking ahead, our future research will focus on addressing challenges within the broader scope of spurious correlations embedded in VLMs.
-This includes using a non-linear probe to distill task-specific core features and mitigating bias.
-Appendix A Appendix / supplemental material
-A.1 Implementation Details
-Datasets overview.
-We describe the dataset details used in our study here:
-The Waterbirds dataset [Sagawa et al., 2019] is a popular benchmark for binary classification tasks, specifically designed to examine spurious correlations.
-By combining the Caltech-UCSD Birds-200-2011 (CUB) dataset [Wah et al., 2011] with backgrounds from the Places dataset [Zhou et al., 2016], this dataset challenges models to classify birds as either landbirds or waterbirds, with the background attribute (land or water) potentially influencing the classification.
-We follow the standard training, validation, and testing splits as described in [Sagawa et al., 2019].
-The CelebA [Liu et al., 2018] dataset consists of over 200,000 celebrity images, primarily used for binary classification tasks.
-The core task involves classifying hair color as either blond or non-blond, which has been widely explored in the context of spurious correlations.
-Interestingly, gender emerges as a spurious attribute in this dataset.
-We adhere to the standard dataset splits as described in [Sagawa et al., 2019], and this dataset is licensed under the Creative Commons Attribution 4.0 International license.
-In addtion to hair color, in Figure 2, we also test the other attributes spuriously correlated with gender.
-See Table 3 for more details.
-Dataset Preprocessing.
-Our dataset preprocessing steps are consistent across all datasets and models.
-Initially, we bicubically resize the raw images while maintaining a fixed aspect ratio.
-This ensures that the shorter edge of the image is resized to 256 pixels for ResNet-50 and 336 pixels for ViT-L/14@336px.
-Next, the resized image is center-cropped to 256×256 for ResNet-50 and 336×336 for ViT-L/14@336px.
-Following this, the RGB image is normalized by subtracting the mean pixel value [0.4815, 0.4578, 0.4082] and dividing by the standard deviation [0.2686, 0.2613, 0.2758], in line with CLIP’s procedure.
-No additional data augmentation is applied after these steps.
-| Attribute Name |
-| Arched Eyebrows | 54932 | 64560 | 39577 | 3701 | ||||||||
-| Attractive | 29920 | 49247 | 64589 | 19014 | ||||||||
-| Bags Under Eyes | 84963 | 44527 | 9546 | 23734 | ||||||||
-| Bald | 94500 | 64557 | 9 | 3704 | ||||||||
-| Bangs | 75612 | 62473 | 18897 | 5788 | ||||||||
-| Big Lips | 65962 | 57595 | 28547 | 10666 | ||||||||
-| Big Nose | 84954 | 39475 | 9555 | 28786 | ||||||||
-| Black Hair | 75725 | 48139 | 18784 | 20122 | ||||||||
-| Blond Hair | 71629 | 66874 | 22880 | 1387 | ||||||||
-| Blurry | 90109 | 64299 | 4400 | 3962 | ||||||||
-| Brown Hair | 71706 | 57872 | 22803 | 10389 | ||||||||
-| Bushy Eyebrows | 87757 | 51627 | 6752 | 16634 | ||||||||
-| Chubby | 93392 | 59989 | 1117 | 8272 | ||||||||
-| Double Chin | 93620 | 61579 | 889 | 6682 | ||||||||
-| Eyeglasses | 92354 | 59895 | 2155 | 8366 | ||||||||
-| Gray Hair | 93563 | 62311 | 946 | 5950 | ||||||||
-| Heavy Makeup | 32157 | 68058 | 62352 | 203 | ||||||||
-| High Cheekbones | 41836 | 47289 | 52673 | 20972 | ||||||||
-| Mouth Slightly Open | 44938 | 39346 | 49571 | 28915 | ||||||||
-| Narrow Eyes | 83877 | 60024 | 10632 | 8237 | ||||||||
-| No Beard | 117 | 26874 | 94392 | 41387 | ||||||||
-| Oval Face | 63330 | 53339 | 31179 | 14922 | ||||||||
-| Pale Skin | 89199 | 66566 | 5310 | 1695 | ||||||||
-| Pointy Nose | 60774 | 57150 | 33735 | 11111 | ||||||||
-| Receding Hairline | 89502 | 60228 | 5007 | 8033 | ||||||||
-| Rosy Cheeks | 84200 | 68045 | 10309 | 216 | ||||||||
-| Smiling | 43688 | 41002 | 50821 | 27259 | ||||||||
-| Straight Hair | 76848 | 51975 | 17661 | 16286 | ||||||||
-| Wavy Hair | 52289 | 58499 | 42220 | 9762 | ||||||||
-| Wearing Earrings | 65206 | 67202 | 29303 | 1059 | ||||||||
-| Wearing Hat | 92112 | 62619 | 2397 | 5642 | ||||||||
-| Wearing Lipstick | 18516 | 67817 | 75993 | 444 | ||||||||
-| Wearing Necklace | 75984 | 67022 | 18525 | 1239 | ||||||||
-| Young | 11167 | 24815 | 83342 | 43446 |
-Model Architecture.
-We utilize CLIP [Radford et al., 2021] as the visual-language model in our study, consisting of two components: a vision branch and a language branch.
-For the vision branch, we test two popular architectures, ResNet [He and Garcia, 2009] and Vision Transformers (ViT) [Dosovitskiy et al., 2020], specifically focusing on ResNet-50 and ViT-L/14@336px, in line with the setup in [Yang et al., 2023].
-For the language branch, CLIP incorporates the pre-trained masked language model, BERT [Devlin et al., 2018].
-Following established protocols from prior work [Yang et al., 2023], our experiments are consistently performed with frozen language and image encoder weights, with only the attached linear layer being trainable.
-Training Details.
-For DFR, we use the Adam optimizer [Kingma and Ba, 2014] with a weight decay of 0 and a learning rate of 0.01.
-The ReduceLROnPlateau scheduler is adopted with a factor of 0.5 and patience of 3.
-The models are trained for 20 epochs with a batch size of 256.
-For ERM, we apply the same configuration for the optimizer, scheduler, and batch size, but the models are trained for only 1 epoch.
-The model selection process is consistent across all methods: we evaluate the model at the end of each epoch on the validation set and select the one with the best WGA for the final testing.
-All accuracy metrics reported in this paper are based on the test set.
-Computational Resources.
-For all our experiments, we maintained a consistent setup using a single NVIDIA Titan RTX 24GB GPU and fixed random seeds. The experiments were conducted using PyTorch 2.0.1+cu117 and Python 3.8.13.
-Evaluation metrics.
-Worst-Group Accuracy (WGA) represents the lowest model accuracy among different groups in the testing set, as defined in Section 3.
-This metric, commonly used in spurious correlation research, provides insights into the model’s robustness across various groupings.
-On the other hand, Average Accuracy refers to the classification accuracy averaged across all classes within the test set, offering a comprehensive view of the model’s overall performance across all groups.
-A.2 Cosine Similarities Distribution on non-spurious correlated prompt image pair
-A.3 PromptCraft workflow
-A.4 Proof of Eq. (4)
-Let be a subspace of and let be a vector in . We denote the closest vector to on by . Let
-be the orthogonal decomposition with respect to .
-By definition lies in , where is the base of subspace and so there exist a vector in such that
-We know that lies in , we thus have
-and so
-Suppose that
-Then
-so by the previous proof. But (the orthogonal decomposition of the zero vector is just ), so , and therefore is in .
-Since the columns of are linearly independent, we have , so , as desired. Let be a vector in and let be a solution of
-Then
-so
-A.5 VisualDistiller Performance on CelebA: Non-blond/blond hair
 | CLIP ViT | CLIP ResNet | ||||||||
@@
 | DFR | no projection | 89.38% | 90.70% | 89.77% | 91.38% | ||||
-A.6 VisualDistiller Performance on CelebA: Other Attributes
 | Attributes |
@@
 | ERM | a corresponding gender image | 94.56% | |||||
-| Brown Hair | DFR | no projection | 95.41% | ||||
-| no projection | 57.27% | ||||||
-| an irrelevant image | 66.36% | ||||||
-| an opposite gender image | 66.36% | ||||||
-| a male and female image | 73.64% | ||||||
-| ERM | a corresponding gender image | 76.36% | |||||
-| Grey Hair | DFR | no projection | 96.36% | ||||
-| no projection | 51.92% | ||||||
-| an irrelevant image | 50.58% | ||||||
-| an opposite gender image | 82.05% | ||||||
-| a male and female image | 87.81% | ||||||
-| ERM | a corresponding gender image | 81.64% | |||||
-| Wavy Hair | DFR | no projection | 83.56% |
-A.7 Zero-shot Classification Performance by Different Prompt
-| WG | Avg | |||||
-| non-blond/blond hair | 99.12% | 97.32% | 10.85% | 26.67% | 10.85% | 85.30% |
-| dark/blond hair | 73.04% | 56.88% | 98.39% | 92.22% | 56.88% | 70.16% |
-| a celebrity with dark/blond hair | 69.59% | 59.75% | 98.59% | 92.78% | 59.75% | 69.85% |
-| a photo of a celebrity with dark/blond hair | 79.58% | 81.91% | 96.33% | 83.33% | 79.58% | 82.92% |
-| A photo of a celebrity with dark/blond hair. | 85.77% | 88.65% | 90.16% | 73.89% | 73.89% | 87.45% |
-| WG | Avg | |||||
-| landbird/waterbird | 98.89% | 85.06% | 24.92% | 52.49% | 24.92% | 87.39% |
-| a photo of a landbird/waterbird | 99.02% | 88.60% | 35.67% | 64.64% | 35.67% | 90.41% |
-| terrestrial/aquatic bird | 97.47% | 69.22% | 34.58% | 74.61% | 34.58% | 90.68% |
-| This is a picture of a landbird/waterbird. | 98.58% | 86.08% | 45.64% | 66.98% | 45.64% | 90.60% |
-A.8 Text Recovered from DRF Linear Probe by Setting Different EOT Index
-| EOT Index | Recovered Text |
-| 6 | tman photo hyuk landbird |
-| 7 | ta mascot lowers !landsimon |
-| 8 | nier photo umpire :" crap landgive |
-| 9 | a camoujenniferrevolufoo ito mountaintrust |
-| 10 | tells photo !! birds ent rattwouldn |
-| 11 | mcr brandon reland !!!!do!stephen give |
-| 12 | tried nfl razzcardinalusa!stupid unk landptv |
-| 13 | lbs photo apocalypse jnr eem !! buucan dotcom land |
-| 14 | hated photo apocalypse !!!:-) !4 bag landarily |
-| 15 | harrison seo thofjuly pals ?! !!! outh rap diamonwonder |
-| 16 | hated photo cyber….there ju $) !!!!navajo land |
-| 17 | nottphoto ! lmaooo !!!landquered |
-| 18 | ghanaian photo scorpio bamcfb -) !rigor!skullgovernors |
-| 19 | fueled photo remy syrian radicals @# !!!!gubernat!!!hedgeesof |
-| 20 | hello photo lulla.) !… !condemn ==>slapped amo! !landcancel |
-| 21 | harris painters afghanistan harris erz flyo…" %; adays arrog.’ soweto popowindothat led window |
-| 22 | lotta pistemails ….reveals .* ; acoladays .… rondo gravy calabhis !allenbird |
-| 23 | cco photo motel horror ninety y …………….showroom ???? !????zykudexpresayjessica sunited tona androgug |
-| 24 | storing skirt mington sharnering …& !.. !!.. !!"… !!!captured callimos |
-| 25 | ello damonsharpe um !!!!!!!! ?!%! automate issboys montgomery mts forms mascokanzzvamosaygoo !aparthegroun |
-| 26 | nels artichoke henry !!.… ………. […] …….. …….. @@: @.… pict !!!!locust child |
-| EOT Index | Recovered Text |
-| 6 | cupcakes dec photo watercrosses |
-| 7 | .@ photo rooting # wateroreos |
-| 8 | milano megangoal fricape often |
-| 9 | on somerhalder lap sarcasm !place scones |
-| 10 | keegan tea supper ?!premiered !!platte fires |
-| 11 | rande clothes brunch !ousness |
-| 12 | jameson pond pepper oh anime cia send !!hermione sergey |
-| 13 | tomas petrovixx ^motherhood " scarf !mainwpwaterbird |
-| 14 | concert riga thypovertls hug mpacific !ce fowjumpers bird |
-| 15 | mug rence poon howling !vie home myra unified !!ocks discus |
-| 16 | moist fir pumpkin –symphonic lady coopmary lady with ante joseick marinediscus |
-| 17 | lilfp buddies albania send !!ilikesandalwood worcester theast !seagull pods |
-| 18 | aro photo of somalia eloqu amazingly davies !!!!!phy swan swan |
-| 19 | a photo friend aldubch !shadesof!!bvb (. duf!!!rice isles swans |
-| 20 | a beer marshmallows !. !twinning duper firstdayofexists ••i !!!!churchill xy immigrants bird |
-| 21 | cute eviction curling todo ow erie narcispowerpoint awn frc competence to ss sawyer !rideelygull bird |
-| 22 | professionally photo bears !!!awesome deep agic blackandwhite !!sl !sabre carlos # gull discus |
-| 23 | a photo fish :( !driven ebo ffl !!.kah !!!salt stockdiscus |
-| 24 | a photo that .? !!!rip !!!!!!!cmo !# !williamspods |
-| 25 | a photo of tko hermes !quot!!sri !mlk !!!with ert cherished mere kiteconcorponies |
-| 26 | a photo with …& !!!!splacchapters pic sizes grin kesh !!wq!!pos !thinswan discus |
+| Brown Hair | DFR | no projection | 95.41% |
+7 Conclusions
+In this study, we explored the capabilities of a CLIP model in manipulating a linear probe from multiple perspectives.
+We showed that the text prompt representations are often tainted by contextual features embedded within the training data.
+In contrast, visual representations demonstrated greater expressiveness, and targeting specific features within these representations proved highly effective for extracting essential information for downstream tasks.
+Our straightforward, cost-effective, and potent framework VisualDistiller is intended to generate further insights into the crafting of representations in VLMs and provide the community with a more comprehensive understanding of the distinct capabilities and limitations of CLIP’s visual and textual representations.
+Limitations and Broader Impacts.
+Mitigating spurious correlations in machine learning models is crucial for developing more reliable and trustworthy AI.
+Regarding privacy and security risks, these are relatively low in our study, as our work builds upon certified VLMs like CLIP.
+Looking ahead, our future research will focus on addressing challenges within the broader scope of spurious correlations embedded in VLMs.
+This includes using a non-linear probe to distill task-specific core features and mitigating bias.

### v3

@@
-Refining Skewed Perceptions in Vision-Language Models through Visual Representations
+Refining Skewed Perceptions in Vision-Language Contrastive Models through Visual Representations
 Abstract
-Large vision-language models (VLMs), such as CLIP, have become foundational, demonstrating remarkable success across a variety of downstream tasks.
+Large vision-language contrastive models (VLCMs), such as CLIP, have become foundational, demonstrating remarkable success across a variety of downstream tasks.
 Despite their advantages, these models, akin to other foundational systems, inherit biases from the disproportionate distribution of real-world data, leading to misconceptions about the actual environment.
-Prevalent datasets like ImageNet are often riddled with non-causal, spurious correlations that can diminish VLM performance in scenarios where these contextual elements are absent.
+Prevalent datasets like ImageNet are often riddled with non-causal, spurious correlations that can diminish VLCM performance in scenarios where these contextual elements are absent.
 This study presents an investigation into how a simple linear probe can effectively distill task-specific core features from CLIP’s embedding for downstream applications.
 Our analysis reveals that the CLIP text representations are often tainted by spurious correlations, inherited in the biased pre-training dataset.
-Empirical evidence suggests that relying on visual representations from CLIP, as opposed to text embedding, is more practical to refine the skewed perceptions in VLMs, emphasizing the superior utility of visual representations in overcoming embedded biases.
-Our codes will be available upon publication.
+Empirical evidence suggests that relying on visual representations from CLIP, as opposed to text embedding, is more effective to refine the skewed perceptions in VLCMs, emphasizing the superior utility of visual representations in overcoming embedded biases.
+Our code can be found here.
 1 Introduction
-Vision-language models (VLMs), a class of multimodal artificial intelligence systems, seamlessly bridge the gap between visual perception and natural language understanding, providing users with a more intuitive way to leverage artificial intelligence for solving daily problems.
+Vision-language contrastive models (VLCMs), a class of multimodal artificial intelligence systems, seamlessly bridge the gap between visual perception and natural language understanding, providing users with a more intuitive way to leverage artificial intelligence for solving daily problems.
 The synergy between visual and linguistic data has significant implications for various applications, including image generation, image captioning, cross-modal retrieval, and visual question answering.
-VLMs like Contrastive Language-Image Pre-training (CLIP) [17] have set new benchmarks across various tasks by contrastively matching semantically closest image and text pairs.
-However, due to the disproportionate distribution embedded in real-world datasets like ImageNet [5] or LAION [19], pre-trained VLMs inherently acquire biases from these large-scale datasets.
+VLCMs like Contrastive Language-Image Pre-training (CLIP) [19] have set new benchmarks across various tasks by contrastively matching semantically closest image and text pairs.
+However, due to the disproportionate distribution embedded in real-world datasets like ImageNet [5] or LAION [21], pre-trained VLCMs inherently acquire biases from these large-scale datasets.
 This phenomenon, known as spurious correlation, refers to patterns that correlate the target class with non-causal contextual elements.
@@
 Similarly, BERT’s [6] peak performance on the argument reasoning comprehension task is largely due to exploiting spurious statistical cues in the dataset, like the negation word “not” [17].
-In this work, we investigate the spurious correlations embedded in foundational VLMs like CLIP and aim to answer the following questions: 1) Does CLIP rely on non-causal “background” features in its decision-making process?
+In this work, we investigate the spurious correlations embedded in foundational VLCMs like CLIP and aim to answer the following questions: 1) Does CLIP rely on non-causal “background” features in its decision-making process?
 If so, how?
@@
 To examine the practicality of zero-shot classification in the presence of spurious correlations, we evaluate the degree of contamination in CLIP’s text representations due to biased pre-training data through extensive statistical analysis.
-Lastly, we show CLIP’s visual representation’s ability to distill core features using the proposed VisualDistiller framework.
-In summary, we make following contribution in this work:
-We show that VLMs like CLIP rely on non-causal spurious features for decision-making, yet linear probing is sufficient to extract key features for various downstream tasks.
+Lastly, we show CLIP’s visual representation’s ability to distill core features using the experimental VisualDistiller framework.
+In summary, in this work, we provide a deeper understanding of the strengths and limitations of CLIP’s visual and textual representations rather than purely benchmarking spurious correlation mitigation methods.
+Our key findings include:
+We show that VLCMs like CLIP rely on non-causal spurious features for decision-making, yet linear probing is sufficient to extract key features for various downstream tasks.
 We find that CLIP’s text embeddings are contaminated by diverse elements, making text embeddings impractical for debiasing the model.
 We demonstrate that using visual embeddings from CLIP to distill visual representations is highly effective.
-The debiased features achieve excellent performance in group accuracy comparable to supervised methods like DFR [11], which offers a more comprehensive understanding of the distinct capabilities and limitations of CLIP’s visual and textual representations.
+The debiased features achieve excellent performance in group accuracy comparable to state-of-the-art supervised methods [11].
+This knowledge can inform the development of fairer and more robust machine learning models, such as unbiased image retrieval algorithms and image generation leveraging CLIP.
 2 Related Work
@@
 This issue contrasts with the training phase, where both groups generally achieve more balanced training accuracy [20, 7].
-[7, 20, 9] highlights that neural networks are prone to a simplicity bias, often emphasizing trivial spurious features while neglecting the essential core features.
+[7, 22, 9] highlight that neural networks are prone to a simplicity bias, often emphasizing trivial spurious features while neglecting the essential core features.
 To address these challenges, substantial research has been dedicated to enhancing robustness against spurious correlations.
@@
 For example, deep feature reweighting (DFR) [11] solves this by retraining the last linear layer of the model (e.g.
-ResNet) by upweighting the underrepresented samples in the loss function, ensuring better handling of minority groups.
+ResNet) by up-weighting the underrepresented samples in the loss function, ensuring better handling of minority groups.
 In scenarios where group labels are unavailable, a common approach involves initially training an auxiliary model using empirical risk minimization (ERM).
@@
 This robust model is typically trained using techniques such as sample balancing [13, 16], or contrastive learning [31, 30, 27] with the inferred group labels.
-Enhancing Group Robustness in VLMs.
-VLMs have gained increased popularity for their ability to perceive the world through multiple modalities.
+Enhancing Group Robustness in VLCMs.
+VLCMs have gained increased popularity for their ability to perceive the world through multiple modalities.
 Previous research has sought to enhance the robustness of vision classifiers by incorporating language features, using techniques such as attention maps [18] and modifications to feature attributes [32].
@@
 Conversely, [27] pioneers a fine-tuning strategy specifically designed to address spurious correlations with group labels in pre-trained multimodal models.
-[3] addresses VLMs’ bias in zero-shot classification by projecting out biased directions in the text embeddings.
-[25] proposed composition-aware hard negatives during training, which improves the model’s ability to understand attributes, relations, and order significantly, leading to better performance on tasks that require compositional understanding.
-Unlike these approaches, our objective is to investigate the inherent skewed perception embedded in all text embeddings (including target class text and spurious attribute text) and explore the possibilities of using visual representations instead to distill the task-specific core features from VLMs like CLIP for downstream tasks, without the need for group annotations.
+[3] addresses VLCMs’ bias in zero-shot classification by projecting out biased directions in the text embeddings.
+[29] proposes composition-aware hard negatives during training, which improves the model’s ability to understand attributes, relations, and order significantly, leading to better performance on tasks that require compositional understanding.
+Unlike these approaches, our objective is to investigate the inherently skewed perception embedded in all text embeddings (including target class text and spurious attribute text) and explore the possibilities of using visual representations instead to distill the task-specific core features from VLCMs like CLIP for downstream tasks, without the need for group annotations.
 3 Preliminaries
@@
 For a given classification task, we have samples , where represents the input features, as the class labels, as the spurious attributes, and as the group labels.
-We examine scenarios of distribution shifts occurring between samples across different groups but from same the same class.
+We examine scenarios of distribution shifts occurring between samples across different groups but from the same class.
 In the Waterbirds dataset [20], we define {landbird, waterbird}, {land background, water background}, and {landbird on land (), landbird on water (), waterbird on land (), waterbird on water ()}.
@@
 The goal of ERM is to minimize the average classification error using a classifier , described mathematically as:
-where is the loss function. To achieve robustness across groups, one aim to minimize the worst-group error:
+where is the loss function. To achieve robustness across groups, one aims to minimize the worst-group error:
 4 A Linear Probe Can Achieve Optimal Task Performance
@@
 For classification tasks in particular, the ideal is to utilize the same representations derived from a pre-trained model across various classification challenges.
-Given the nature of spurious correlations — where a feature deemed spurious for one task may be essential for another — we expect the VLMs to capture a broad spectrum of nuanced visual information, and removing the spurious feature by specialized modules.
-This section delves into the presence of spurious correlations within VLMs and explores whether a simple linear probe can deliver optimal performance.
-4.1 Unraveling Spurious Correlations in VLMs
-In this section, we investigate the existence of spurious features within CLIP through a comparative analysis of zero-shot binary classification performance on the Waterbirds dataset.
-We use the text encoder in CLIP to obtain representations for the text labels (“a photo of a landbird” and “a photo of a waterbird”) and the image encoder to extract features from the images.
-The most similar label, which the model would assign to the image later, is determined by calculating the cosine similarity between the image and text features.
-To show that spurious correlation impairs the performance of CLIPs, we conduct two experiments: zero-shot classification using the original Waterbirds dataset (with natural background) and using a modified version of the Waterbirds dataset from which the background have been erased based on mask (available in Waterbird dataset).
-Figure 1 shows one example of the two version of Waterbird dataset.
-Figure 2 presents the group accuracy changes across these two scenarios.
-In the first scenario (using original data), the models exhibit uneven accuracies across majority and minority groups, reflecting the unbalanced group robustness across the dataset.
-When the backgrounds are removed, the accuracy in recognizing and (minority groups) boosts, as the confounding elements are no longer able to mislead the model.
-Despite the relatively small change on majority , we see consistent accuracy drop in majority , across three different architectures.
-This indicates that the model depends more heavily on background information when classifying waterbird-on-water samples compared to landbird-on-land samples.
-Conclusion 1.
-The disparity in performance and drastic accuracy change in both minority and majority group, confirms our hypothesis that visual representations in current models are entangled with spurious features that significantly impair classification performance.
-This raises an fundamental question: does this imply that we are unable to achieve flawless task execution on CLIP representations when faced with spurious features?
-If not, how?
+Given the nature of spurious correlations — where a feature deemed spurious for one task may be essential for another — we expect the VLCMs to capture a broad spectrum of nuanced visual information and remove the spurious feature by specialized modules.
+This section delves into the presence of spurious correlations within VLCMs and explores whether a simple linear probe can deliver optimal performance.
+4.1 Unraveling Spurious Correlations in VLCMs
+We first examine the presence of spurious correlations in CLIP by evaluating zero-shot binary classification performance on the Waterbirds dataset.
+Using CLIP’s text and image encoders, we compute the cosine similarity between image features and textual labels (“a photo of a landbird” and “a photo of a waterbird”).
+To assess the impact of spurious correlations, we compare classification performance on the original dataset (with natural backgrounds) and a modified version where backgrounds are removed using segmentation masks (Figure 1).
+As shown in Figure 2(a), accuracy disparities emerge between majority and minority groups when using the original dataset.
+Removing backgrounds improves the classification accuracy of minority groups (, ), suggesting that background cues previously misled the model.
+Meanwhile, accuracy drops for the majority group , indicating a stronger reliance on background features for waterbird-on-water samples, aligning with prior work findings [27, 26].
+These results reinforce that CLIP representations encode spurious features, which can hinder classification performance.
+This raises a key question: can CLIP-based models achieve robust classification despite these spurious correlations?
 4.2 Assessing the Expressiveness of CLIP’s Visual Representations under Linear Probing
 In order to see the upper limit of CLIP visual representation with linear transformation, we applied DFR [11], a state-of-the-art method that retrains the last linear layer of the image classifiers (e.g.
-ResNet) by upweighting the underrepresented samples in the loss function, to 29 attribute classification challenges on the CelebA dataset, where each attribute demonstrated a gender-biased distribution.
+ResNet) by up-weighting the underrepresented samples in the loss function, to Waterbirds dataset.
+Following DFR’s implementation, a linear layer is attached to the CLIP image encoder to facilitate binary classification, with updates restricted solely to the weights of the linear layer.
+As a supervised method (knowing the group label of each sample), DFR usually signifies the peak performance that a linear layer can attain and reduces the impact of spurious attributes without altering the primary network.
+High accuracy in these groups suggests that the standard CLIP image encoder successfully captures essential task-related features.
+Figure 2(b) demonstrates the peak classification performance across Waterbirds groups with a CLIP image encoder and a linear projection head.
+Under DFR, the worst-group accuracy (WGA) improves to 82% (ViT-B/32) and 80% (ResNet), exceeding the performance of CLIP model fine-tuned for mitigating spurious correlation [27].
+This highlights that CLIP representations effectively encode the essential information needed for the original classification task.
+To further validate the richness of CLIP’s visual representation, in Figure 3, we applied DFR to 29 attribute classification challenges on the CelebA dataset, where each attribute demonstrated a gender-biased distribution.
 The size of each group w.r.t.
-each attributes can be found in supplementary material.
+each attribute can be found in supplementary material.
 Likewise, they also involve four groups based on gender and attribute presence: female without [attribute] (), male without [attribute] (), female with [attribute] (), and male with [attribute] ().
-Following DFR’s implementation, a linear layer is attached to the CLIP image encoder to facilitate binary classification, with updates restricted solely to the weights of the linear layer.
-As a supervised method (knowing the group label of each sample), DFR usually signifies the peak performance that a linear layer can attain and reduce the impact of spurious attributes without altering the primary network.
-High accuracy in these groups suggests that the standard CLIP image encoder successfully captures essential task-related features, not merely relying on these features for predictions under ERM.
 Figure 3 showcases the outcomes for various binary attribute classifications with a CLIP ViT-L/14 image encoder and a linear projection head, sorted by ascending average group accuracy (a weighted average determined by the size of each group).
 The spectrum of attributes ranged from subtle features like straight hair and narrow eyes to more overt characteristics such as eyeglasses, baldness, and hats.
-Notably, the DFR approach strategy enabled a majority of the attributes — over 25 out of 29 — to achieve more than 75% worst group accuracy (WGA), with more than 23 attribute classification tasks surpassing 80% average group accuracy.
+Notably, the DFR approach strategy enabled a majority of the attributes — over 25 out of 29 — to achieve more than 75% WGA, with more than 23 attribute classification tasks surpassing 80% average group accuracy, and even outperforming calibrated fine-tuned CLIP [28].
 The five attributes with the highest accuracies exceeded 95% in both the worst and average group accuracy measures.
 These results underscore how fine-grained the CLIP’s visual representations are, capturing a comprehensive spectrum of visual information, including subtle features that are typically challenging for human perception.
-Conclusion 2.
+Finding 1.
 Visual representations learned by CLIP are adept at extracting nuanced features within images for various tasks by linear transformation.
 5 Language Representations are More Noisy than Expected
-VLMs like CLIP prevail partly because of their capability to perform zero-shot inferences guided by intuitive language cues.
+VLCMs like CLIP prevail partly because of their capability to perform zero-shot inferences guided by intuitive language cues.
 As demonstrated in the previous section, techniques such as DFR guide us toward identifying an optimal linear probe that can effectively discern variations in core features.
@@
 This section explores the biases inherent in zero-shot classification prompts.
-VLMs are trained to align the representations of images with their corresponding captions via cosine similarities.
+VLCMs are trained to align the representations of images with their corresponding captions via cosine similarities.
 Ideally, one might expect the representation of “a photo of a dog” to solely encapsulate the dog’s key features without incorporating ambient elements like lawns.
@@
 This pattern is consistent across various tested pairs, underscoring the substantial presence of context-related features in CLIP text representations that are not explicitly present in the prompts.
-For the prompt and image pair with less pronounced correlations, like “a photo of a dog” to forest and desert, we did not observe the same level of disparity in mean and median value, see supplementary materials for more details.
-Conclusion 3.
+For the prompt and image pair with less pronounced correlations, like “a photo of a dog” to forest and desert, we did not observe the same level of disparity in mean and median values.
+For disparity on additional text and image pairs, see supplementary materials.
+Finding 2.
 Using text representations for zero-shot classification or debiasing with the representations from spurious attribute prompts [3] could lead to unexpected outcomes, due to the embedded non-target features.
-6 Visual Representations Can Refine the Skewed Perception in VLMs
+6 Visual Representations Can Refine the Skewed Perception in VLCMs
 Due to the noisy nature of the text embeddings from CLIP, we wonder if we can use images to construct an optimal linear probe.
 The availability of background images in the Waterbirds dataset inspires us to explore whether we can use the background images to remove non-core features from the representations of the original Waterbirds images.
-In Figure 6, we demonstrate our proposed framework VisualDistiller, which achieves high group accuracy with only ERM and a simple projection.
-We opted for an ERM linear probe (trained by only one epoch, not knowing the group labels of training samples) over zero-shot text classification due to the existence of context-related features in text representations, which compromise classification reliability.
+In Figure 6, we demonstrate an experimental framework VisualDistiller, using only ERM along with a few inference and projection steps —making it significantly more lightweight than retraining the entire CLIP framework.
+We opted for an ERM linear probe (trained by only one epoch, without knowing the group labels of training samples) over zero-shot text classification due to the existence of context-related features in text representations, which compromise classification reliability.
 The process is as follows: After encoding the target image via CLIP, we obtained the target image feature .
@@
 is then processed through the ERM-trained linear probe to produce the final classification result.
-For simplicity and ease of implementation, we assume Euclidean orthogonality, as using non-Euclidean orthogonality—specifically, employing the inverse of the covariance matrix in the neighborhood space as the metric for inner product calculations—did not yield significant improvement.
-The definition of “background” image varies with the dataset.
-For the Waterbirds dataset, artificially created with images from Places [29] and CUB [23], we defined a range of “background” conditions from least related (random images from Places) to most related (natural environments like lakes and forests) to specific backgrounds used in Waterbirds.
-Figure 7 shows examples of background images with varying levels of semantic similarity to the Waterbird sample image’s background, which is used in Table 1.
-In Table 1, we demonstrate that the experiments using semantically more similar “background” images yields higher WGA in Waterbirds.
+For simplicity and ease of implementation, we assume Euclidean orthogonality, proven effective in [12, 15].
+Using non-Euclidean orthogonality — specifically, employing the inverse of the covariance matrix in the neighborhood space as the metric for inner product calculations — did not yield substantial improvement.
+The definition of a “background” image varies with the dataset.
+For the Waterbirds dataset, artificially created with images from Places [33] and CUB [25], we defined a range of “background” conditions from loosely related (retrieved111Here, “retrieved” refers to the process of encoding all images in the Places dataset using the CLIP ViT-L/14 image encoder and then identifying the top 50 images whose embeddings are most similar to the text embedding using K-nearest neighbors.
+images from Places using “a photo of indoor or outdoor background”) to more strictly related (retrieved images from Places using “a photo of nature background”) to corresponding backgrounds used in Waterbirds.
+Figure 7 shows examples of background images with varying levels of semantic similarity, which is used in Table 1.
+In Table 1, we demonstrate that the experiments using semantically more similar “background” images yield higher WGA in Waterbirds.
 Besides, transitioning to an ERM-trained linear probe enhances WGA further by focusing more sharply on core features, unlike the “contaminated” text representations from CLIP.
@@
 The VisualDistiller can achieve the WGA of 82.40% without knowing the background image category (20 random images from nature, ViT), only a few points from supervised DFR’s 85.67% performance.
+We also observe a minor decrease in the average accuracy of our proposed experiments compared to ERM without projection on the original Waterbirds dataset.
+This is expected, and widely exists in similar framework [27, 28], as ERM prioritizes overall accuracy at the expense of minority group performance — the main optimization subject of the work.
 | CLIP ViT | CLIP ResNet | |||||||||||
-WG | Avg | WG | Avg | ||||||
+WGA | Avg | WGA | Avg | ||||||
 | no projection | n/a | 35.67% | 90.41% | 36.14% | 92.89% | |||||||
@@
 | corresponding subclass text | 1 | 46.57% | 89.43% | 44.24% | 93.09% | |||||||
-| a random image from Places | 1 | 42.68% | 90.56% | 48.29% | 90.65% | |||||||
-| a random image from same class | 1 | 54.83% | 86.95% | 66.82% | 86.41% | |||||||
-| a random image from same subclass | 1 | 57.94% | 87.51% | 71.34% | 81.68% | |||||||
+| an image retrieved by “a photo of indoor or outdoor background” | 1 | 42.68% | 90.56% | 48.29% | 90.65% | |||||||
+| an image retrieved by corresponding class text | 1 | 54.83% | 86.95% | 66.82% | 86.41% | |||||||
+| an image retrieved by corresponding subclass text | 1 | 57.94% | 87.51% | 71.34% | 81.68% | |||||||
 | Zero-shot | the corresponding background | 1 | 55.45% | 87.55% | 75.23% | 87.65% | ||||||
@@
 | 3 | 70.09% | 96.31% | 62.15% | 94.71% | ||||||||
-| random images from Places | 10 | 71.81% | 96.06% | 63.08% | 94.08% | |||||||
+| images retrieved by “a photo of indoor or outdoor background” | 10 | 71.81% | 96.06% | 63.08% | 94.08% | |||||||
 | 1 | 77.73% | 97.33% | 62.93% | 95.61% | ||||||||
@@
 | 10 | 81.46% | 96.26% | 61.53% | 91.17% | ||||||||
-| random images from nature | 20 | 82.40% | 95.03% | 62.77% | 90.15% | |||||||
+| images retrieved by “a photo of nature background” | 20 | 82.40% | 95.03% | 62.77% | 90.15% | |||||||
 | 1 | 81.93% | 96.20% | 73.52% | 93.60% | ||||||||
 | 3 | 86.29% | 95.47% | 78.82% | 91.74% | ||||||||
-| random images from same class | 10 | 87.07% | 93.45% | 73.99% | 89.86% | |||||||
+| images retrieved by corresponding class text | 10 | 87.07% | 93.45% | 73.99% | 89.86% | |||||||
 | 1 | 84.27% | 95.84% | 74.30% | 94.09% | ||||||||
 | 3 | 87.54% | 94.15% | 79.75% | 92.05% | ||||||||
-| random images from same subclass | 10 | 87.85% | 93.35% | 72.90% | 89.82% | |||||||
+| images retrieved by corresponding subclass text | 10 | 87.85% | 93.35% | 72.90% | 89.82% | |||||||
 | corresponding background | n/a | 88.16% | 96.71% | 79.28% | 93.83% | |||||||
@@
 Additionally, we applied VisualDistiller to the CelebA dataset classifying if the celebrity’s hair color is blond.
-Here, we used images of celebrities without hair as “background” vectors source to mitigate non-hair related features.
-For examples of “background” image in this experiment, see Figure 8.
+Here, we used retrieved222Here, “retrieved” refers to the process of encoding all images in the CelebA dataset using the CLIP ViT-L/14 image encoder and then identifying the top 50 images whose embeddings are most similar to the query embedding using K-nearest neighbors.
+images of celebrities without hair as “background” vector source to mitigate non-hair related features.
+For examples of “background” images in this experiment, see Figure 8.
 Despite real-world limitations preventing the exact matching of these “background” conditions, using a set of bald celebrity images proved effective.
@@
 More experiments on other CelebA attributes can be found in Table 3.
-Note that the proposal of VisualDistiller purely aim to validate the effectiveness of visual represention over the text representation, hence we do not seek to benchmark with other methods like supervised DFR.
+Note that the proposal of VisualDistiller purely aims to validate the effectiveness of visual representation over the text representation, hence we do not seek to benchmark with other methods like supervised DFR.
 | CLIP ViT | CLIP ResNet | ||||||||
-WG | Avg | WG | Avg | ||||
+WGA | Avg | WGA | Avg | ||||
 | no projection | 47.22% | 94.78% | 38.89% | 95.29% | |||||
@@
 | an irrelevant image | 58.89% | 93.81% | 55.56% | 94.38% | |||||
-| an opposite gender image | 66.67% | 85.45% | 66.11% | 87.98% | |||||
-| a male and female image | 79.37% | 86.21% | 81.11% | 87.43% | |||||
-| ERM | a corresponding gender image | 83.88% | 87.60% | 83.33% | 87.76% | ||||
+| an image retrieved by opposite gender text | 66.67% | 85.45% | 66.11% | 87.98% | |||||
+| two images retrieved by corresponding and opposite gender text | 79.37% | 86.21% | 81.11% | 87.43% | |||||
+| ERM | an image retrieved by corresponding gender text | 83.88% | 87.60% | 83.33% | 87.76% | ||||
 | DFR | no projection | 89.38% | 90.70% | 89.77% | 91.38% | ||||
 | Attributes |
-Minority Group Accuracy | ||||
+Worst Group Accuracy | ||||
 | no projection | 88.24% | ||||||
 | an irrelevant image | 94.61% | ||||||
-| an opposite gender image | 95.97% | ||||||
+| an image retrieved by opposite gender text | 95.97% | ||||||
 | a male and female image | 96.97% | ||||||
-| ERM | a corresponding gender image | 94.85% | |||||
+| ERM | an image retrieved by corresponding gender text | 94.85% | |||||
 | Black Hair | DFR | no projection | 95.48% | ||||
@@
 | an irrelevant image | 70.44% | ||||||
-| an opposite gender image | 97.97% | ||||||
+| an image retrieved by opposite gender text | 97.97% | ||||||
 | a male and female image | 98.29% | ||||||
-| ERM | a corresponding gender image | 94.56% | |||||
+| ERM | an image retrieved by corresponding gender text | 94.56% | |||||
 | Brown Hair | DFR | no projection | 95.41% |
@@
 In contrast, visual representations demonstrated greater expressiveness, and targeting specific features within these representations proved highly effective for extracting essential information for downstream tasks.
-Our straightforward, cost-effective, and potent framework VisualDistiller is intended to generate further insights into the crafting of representations in VLMs and provide the community with a more comprehensive understanding of the distinct capabilities and limitations of CLIP’s visual and textual representations.
-Limitations and Broader Impacts.
-Mitigating spurious correlations in machine learning models is crucial for developing more reliable and trustworthy AI.
-Regarding privacy and security risks, these are relatively low in our study, as our work builds upon certified VLMs like CLIP.
-Looking ahead, our future research will focus on addressing challenges within the broader scope of spurious correlations embedded in VLMs.
-This includes using a non-linear probe to distill task-specific core features and mitigating bias.
+Our straightforward, cost-effective, and potent framework VisualDistiller is intended to generate further insights into the crafting of representations in VLCMs and provide the community with a more comprehensive understanding of the distinct capabilities and limitations of CLIP’s visual and textual representations.
+Broader Impacts.
+Mitigating spurious correlations in machine learning models is essential for building reliable and trustworthy AI.
+For instance, recognizing that text embeddings contain spurious biases helps us approach their use in image retrieval with caution.
+In future work, we aim to enhance fairness in image retrieval by leveraging distilled image embeddings as a surrogate for query embeddings, reducing bias in retrieved results.